### v1

# Tidal perturbations of an extreme mass ratio inspiral around a Kerr black hole

Tidal perturbations of an extreme mass ratio inspiral around a Kerr black hole
[…]
Dipartimento di Fisica e Geologia, Università di Perugia, I.N.F.N. Sezione di Perugia,
Via Pascoli, I-06123 Perugia, Italy
Center of Gravity, Niels Bohr Institute, Copenhagen University,
Blegdamsvej 17, DK-2100 Copenhagen Ø, Denmark
Nordita, KTH Royal Institute of Technology and Stockholm University,
Hannes Alfvéns väg 12, SE-106 91 Stockholm, Sweden
§ William H. Miller III Department of Physics and Astronomy, Johns Hopkins University, 3400 North Charles Street, Baltimore, Maryland, 21218, USA
Max Planck Institute for Gravitational Physics (Albert Einstein Institute), Am Mühlenberg 1, Potsdam 14476, Germany
We determine the metric of a Kerr black hole subject to external tidal fields using metric reconstruction techniques. Working within the Newman–Penrose formalism, we solve the Teukolsky master equation for static, quadrupolar modes associated with a slowly varying tidal environment, and reconstruct the corresponding metric perturbation in the outgoing radiation gauge. As an application, we derive the secular Hamiltonian governing the motion of a test particle in the tidally deformed Kerr spacetime and investigate long-term tidal effects relevant to extreme-mass-ratio inspirals. In particular, we compute tidal-induced shifts of the innermost stable circular orbit and the light ring. We find that these tidal corrections are strongly spin dependent, with significantly larger effects for retrograde orbits around rapidly rotating black holes. Our results provide a fully analytic framework for studying tidal interactions and secular dynamics in rotating black-hole spacetimes, with direct applications to gravitational-wave modeling and tests of gravity in the strong-field regime.
1 Introduction
A central goal in black-hole perturbation theory is to obtain a tractable description of gravitational, electromagnetic, and neutrino perturbations of the Kerr spacetime. The pioneering work of Teukolsky [Teukolsky:1972my, Teukolsky:1973ha], based on the Newman–Penrose formalism [Newman:1961qr], established separable wave equations for the Weyl scalars $\psi_{0}$ and $\psi_{4}$, which encode the behavior of Kerr metric perturbations in vacuum. Together with earlier developments in Kerr geometry [Carter:1968rr, Chandrasekhar:1985kt], these results laid the mathematical foundations for the study of black-hole stability, tidal interactions, and superradiant phenomena [Press:1972zz, Teukolsky:1974yv].
Working with Weyl scalars, rather than the metric perturbation itself, provides direct access to physically relevant observables such as energy and angular momentum fluxes [Chrzanowski:1975wv]. This framework includes prescriptions for quantifying the absorption of mass and angular momentum by a black hole exposed to external radiation [Poisson:2004cw]. The Teukolsky equation also enables studies of quasinormal modes and their couplings [Leaver:1985ax, Green:2022htq]. More broadly, it has been widely applied to problems ranging from the stability of the Kerr spacetime to waveform modeling, including hybrid approaches that combine black-hole perturbation theory with post-Newtonian expansions [Damour:2009wj], as well as investigations of horizon absorption and tidal torquing in Kerr spacetimes [Chatziioannou:2012gq].
However, some applications require explicit knowledge of the metric perturbation rather than the Weyl scalars alone. This is the case for the computation of gravitational self-force effects, which depend on local properties of the spacetime geometry [vandeMeent:2015lxa], for second-order self-force calculations, which require the metric perturbation at first order as an input [Pound:2012nt], and more generally for extensions of black-hole perturbation theory beyond the linear regime [Loutrel:2020wbw, Ripley:2020xby].
Metric reconstruction techniques based on Hertz or Debye potentials [PhysRevLett.41.203, Kegeles:1979an] allow one to recover the spacetime perturbation from the Weyl scalars. These methods are essential for the calculation of Geroch–Hansen multipole moments [Geroch:1973am, LeTiec:2020bos], the construction of initial data for binaries of rotating black holes, and the development of waveform models [Yunes:2005ve].
One important application concerns the tidal deformation of black holes, which in the linear regime is described by vacuum perturbations of the Kerr metric. Previous studies address fundamental questions related to the relativistic theory of tidal Love numbers, which vanish for Schwarzschild black holes but exhibit a richer structure in the spinning case, including dissipative contributions associated with horizon absorption [Binnington:2009bb, Hartle:1973zz, Poisson:2004cw, Chatziioannou:2012gq, Chia:2020yla, LeTiec:2020bos, OSullivan:2014ywd]. The nonvanishing dissipative tidal response of Kerr black holes enters directly into the modeling of gravitational-wave signals from compact binaries: it contributes to the phase evolution and waveform amplitude, thereby affecting both parameter estimation and tests of the nature of compact objects [Tagoshi:1997jy, Alvi:2001mx, Yunes:2009ef, Yunes:2010zj, Flanagan:2007ix, Hinderer:2007mb, Baiotti:2010xh, Chatziioannou:2020pqz, Cardoso:2017cfl, Cardoso:2021qqu]. Beyond astrophysical applications, tidal perturbations also provide insight into the geometry and dynamics of black holes, for example through changes in the horizon area [Poisson:2009qj], couplings between spin and tidal fields [Poisson:2014gka], and modifications of the geodesic structure of tidally deformed spacetimes [Pani:2015hfa, Cardoso:2021qqu].
In this paper, we apply metric reconstruction techniques to determine the spacetime metric of a tidally deformed rotating black hole, up to quadrupole order in the tidal deformation. Specifically, we consider a Kerr black hole of mass $M$ and angular momentum $J$, whose world line $\mathscr{L}$ lies in an external spacetime influenced by tidal fields. We assume that these tidal fields vary slowly in space and time.
We are interested in deriving the metric in the near-zone region, defined as the region at distances $r$ from the world line such that $r\ll\mathscr{R}$, where $\mathscr{R}$ denotes the tidal length scale of the external spacetime. This requirement imposes the key condition that the size of the black hole, characterized by its mass $M$, must be small compared to the tidal scale $\mathscr{R}$, i.e., $M\ll\mathscr{R}$. This hierarchy ensures that the tidal deformation can be treated perturbatively and that, on scales larger than $\mathscr{R}$, the external spacetime is not significantly affected by the presence of the black hole [Poisson:2009qj].
We apply our result for the metric perturbation of a Kerr black hole to study the secular dynamics of a test particle in the tidally deformed spacetime. This application is particularly relevant for Extreme-Mass-Ratio Inspiral (EMRI) systems, which consist of a stellar-mass compact object inspiraling into a supermassive black hole with a mass ratio of order $10^{-5}$$10^{-6}$ or smaller [Berry:2019wgg, Barack:2018yly]. EMRIs can emit gravitational-wave signals lasting up to $\mathscr{O}(10^{5})$ orbital cycles in the millihertz LISA band [Barack:2018yly], making them natural laboratories for probing secular phenomena and prime targets for space-based detectors such as LISA.
We use the secular Hamiltonian obtained from the reconstructed metric perturbation to investigate the properties of two characteristic orbits for massive and massless particles in a tidally deformed spacetime, namely the Innermost Stable Circular Orbit (ISCO) and the light ring (LR), respectively. In particular, we compute the tidal-field-induced shifts in the location and properties of these two orbits, and analyze how these shifts depend on the spin of the rotating black hole. We find that both the ISCO and the LR experience spin-dependent tidal shifts whose magnitude and sign depend on the relative orientation between the orbital angular momentum and the black-hole spin. In particular, for a rapidly rotating black hole, retrograde orbits exhibit substantially larger tidal corrections than prograde ones. [^1] We refer to orbits with angular momentum aligned (anti-aligned) with the black-hole spin as prograde (retrograde). In the literature, these are sometimes called co-rotating and counter-rotating orbits.
The paper is organized as follows. In Sec. 2, we introduce the formalism for describing tidal perturbations of a Kerr black hole and present the Teukolsky master equation in the Hartle–Hawking tetrad, providing a general solution for static modes, which we then specialize to the case of slowly varying tidal fields. In Sec. 3, we summarize the metric reconstruction procedure, and in Sec. 4, we apply it to the Kerr spacetime, providing explicit analytic expressions for the components of the metric perturbation at quadrupole order. In Sec. 5, we describe the procedure for computing the secular dynamics of a test particle in a tidally deformed spacetime and derive the secular Hamiltonian perturbatively to first order in the tidal fields, using the reconstructed metric obtained in the previous section. We then use these results in Sec. 6 to study how the properties of the ISCO and the LR are affected by tidal fields. Finally, Sec. 7 summarizes our conclusions and outlines possible directions for future work.
Throughout this paper, we employ geometrized units with $G=c=1$ and adopt the “mostly plus” metric signature (+++)(-+++). Greek indices run over spacetime coordinates, while Latin indices run over spatial coordinates.
2 Teukolsky master equation
To compute perturbations of the Kerr metric induced by external tidal fields, we employ the Teukolsky Master Equation (TME), first introduced by Teukolsky in his pioneering works [Teukolsky:1973ha, Teukolsky:1974yv]. The TME provides a powerful framework for analyzing linear perturbations of rotating black holes in terms of gauge-invariant curvature variables.
We start by considering a linear perturbation of the Kerr metric,
$g_{\mu\nu}=\tilde{g}_{\mu\nu}+h_{\mu\nu},$ (2.1)
where $\tilde{g}_{\mu\nu}$ denotes the metric of the unperturbed Kerr black hole, and $h_{\mu\nu}$ represents the tidal perturbation, which we compute explicitly up to quadrupole order.
We will work in ingoing Kerr coordinates $(v,r,z,\psi)$. These coordinates are related to the usual Boyer–Lindquist coordinates $(t,r,\theta,\phi)$ through the transformation
$v=t+r_{*},\qquad\psi=\phi+r_{\sharp},\qquad z=\cos\theta,$ (2.2)
where $r_{*}$ and $r_{\sharp}$ are defined as
$r_{*}=r+\frac{2Mr_{+}}{r_{+}-r_{-}}\ln\!\left(\frac{r-r_{+}}{2M}\right)-\frac{2Mr_{-}}{r_{+}-r_{-}}\ln\!\left(\frac{r-r_{-}}{2M}\right),$ (2.3)
$r_{\sharp}=\frac{a}{r_{+}-r_{-}}\ln\!\left(\frac{r-r_{+}}{r-r_{-}}\right).$ (2.4)
Here $r_{\pm}=M\pm\sqrt{M^{2}-a^{2}}$ denote the outer and inner horizons of the Kerr spacetime, and $a=J/M$ is the spin parameter. In contrast to Boyer–Lindquist coordinates, ingoing Kerr coordinates remain regular at the event horizon. In these coordinates, the Kerr metric takes the form
$ds^{2}={}$ $-\left(1-\frac{2Mr}{\Sigma}\right)dv^{2}+2\,dv\,dr-2a(1-z^{2})\left(\frac{2Mr}{\Sigma}\,dv+dr\right)d\psi$
$+\frac{\Sigma}{1-z^{2}}\,dz^{2}+\frac{1-z^{2}}{\Sigma}\left[(r^{2}+a^{2})^{2}-a^{2}\Delta(1-z^{2})\right]d\psi^{2},$ (2.5)
where $\Sigma=r^{2}+a^{2}z^{2}=\zeta\bar{\zeta}$, with $\zeta=r-iaz$, and $\Delta=r^{2}-2Mr+a^{2}$.
Teukolsky’s key insight was to reformulate the linearized Einstein equations in terms of curvature perturbations rather than metric perturbations. Specifically, starting from the Weyl tensor $C_{\mu\nu\rho\sigma}$ and writing the metric as in Eq. (2.1), one can extract the part of the Weyl tensor that depends linearly on the perturbation $h_{\mu\nu}$. Following Ref. [Berens:2024czo], we denote this contribution by $C_{\mu\nu\rho\sigma}^{(1)}$. Using this quantity, one can define two complex scalars, $\psi_{0}$ and $\psi_{4}$, which encode the two independent degrees of freedom of the metric perturbation $h_{\mu\nu}$. They are defined as
$\psi_{0}$ $\equiv C_{\mu\nu\rho\sigma}^{(1)}\,l^{\mu}m^{\nu}l^{\rho}m^{\sigma},$ (2.6a)
$\psi_{4}$ $\equiv C_{\mu\nu\rho\sigma}^{(1)}\,n^{\mu}\bar{m}^{\nu}n^{\rho}\bar{m}^{\sigma},$ (2.6b)
where $(l^{\mu},n^{\mu},m^{\mu},\bar{m}^{\mu})$ is a Newman–Penrose null tetrad satisfying the normalization and orthogonality conditions
$l^{\mu}l_{\mu}=n^{\mu}n_{\mu}=m^{\mu}m_{\mu}=\bar{m}^{\mu}\bar{m}_{\mu}=l^{\mu}m_{\mu}=l^{\mu}\bar{m}_{\mu}=n^{\mu}m_{\mu}=n^{\mu}\bar{m}_{\mu}=0,$ (2.7a)
$-l^{\mu}n_{\mu}=m^{\mu}\bar{m}_{\mu}=1.$ (2.7b)
A common choice for the Newman–Penrose tetrad is the Kinnersley tetrad $\{l^{\mu}_{\rm K},n^{\mu}_{\rm K},m^{\mu}_{\rm K},\bar{m}^{\mu}_{\rm K}\}$ [Kinnersley:1969zza], which is aligned with the principal null directions of the Kerr spacetime. In ingoing Kerr coordinates $(v,r,z,\psi)$, it takes the form [Yunes:2005ve]
$l^{\mu}_{\rm K}$ $=\left(2\frac{r^{2}+a^{2}}{\Delta},\,1,\,0,\,\frac{2a}{\Delta}\right),$ (2.8a)
$n^{\mu}_{\rm K}$ $=\frac{1}{2\Sigma}\left(0,\,-\Delta,\,0,\,0\right),$ (2.8b)
$m^{\mu}_{\rm K}$ $=\frac{\sqrt{1-z^{2}}}{\sqrt{2}\,\bar{\zeta}}\left(ia,\,0,\,-1,\,\frac{i}{1-z^{2}}\right),$ (2.8c)
$\bar{m}^{\mu}_{\rm K}$ $=\frac{\sqrt{1-z^{2}}}{\sqrt{2}\,\zeta}\left(-ia,\,0,\,-1,\,-\frac{i}{1-z^{2}}\right).$ (2.8d)
Even though the use of ingoing Kerr coordinates ensures the regularity of the metric at the future event horizon, this does not guarantee that the Weyl scalars $\psi_{0}$ and $\psi_{4}$ remain regular there. Regularity of the Weyl scalars can instead be achieved by adopting the Hartle–Hawking (HH) tetrad $\{l^{\mu}_{\rm HH},n^{\mu}_{\rm HH},m^{\mu}_{\rm HH},\bar{m}^{\mu}_{\rm HH}\}$ [Hawking:1972hy]. The null vectors $l^{\mu}_{\rm K}$ and $n^{\mu}_{\rm K}$ of the Kinnersley tetrad are related to those of the HH tetrad, $l^{\mu}_{\rm HH}$ and $n^{\mu}_{\rm HH}$, through the transformations
$l^{\mu}_{\rm HH}=\frac{\Delta}{2(r^{2}+a^{2})}\,l^{\mu}_{\rm K},\qquad n^{\mu}_{\rm HH}=\frac{2(r^{2}+a^{2})}{\Delta}\,n^{\mu}_{\rm K},$ (2.9)
while the complex null vector $m^{\mu}$ is the same in both tetrads [Poisson:2004cw, LeTiec:2020bos]. Explicitly, the HH tetrad takes the form
$l^{\mu}_{\rm HH}$ $=\left(1,\,\frac{\Delta}{2(r^{2}+a^{2})},\,0,\,\frac{a}{r^{2}+a^{2}}\right),$ (2.10a)
$n^{\mu}_{\rm HH}$ $=\left(0,\,-\frac{r^{2}+a^{2}}{\Sigma},\,0,\,0\right),$ (2.10b)
$m^{\mu}_{\rm HH}$ $=\frac{\sqrt{1-z^{2}}}{\sqrt{2}\,\overline{\zeta}}\left(ia,\,0,\,-1,\,\frac{i}{1-z^{2}}\right),$ (2.10c)
$\bar{m}^{\mu}_{\rm HH}$ $=\frac{\sqrt{1-z^{2}}}{\sqrt{2}\,\zeta}\left(-ia,\,0,\,-1,\,-\frac{i}{1-z^{2}}\right).$ (2.10d)
Moreover, the Weyl scalars in the HH tetrad are related to those in the Kinnersley tetrad by [Chandrasekhar:1985kt]
$\psi_{0}^{\rm HH}=\frac{\Delta^{2}}{4(r^{2}+a^{2})^{2}}\,\psi_{0}^{\rm K},\qquad\psi_{4}^{\rm HH}=\frac{4(r^{2}+a^{2})^{2}}{\Delta^{2}}\,\psi_{4}^{\rm K}.$ (2.11)
Since we work exclusively with the HH tetrad in what follows, we will omit the superscript “HH” from the Weyl scalars unless otherwise stated.
In ingoing Kerr coordinates $(v,r,z,\psi)$ and in the HH tetrad, the TME reads
$\begin{split}\Bigg(&-a^{2}(1-z^{2})\,\partial_{v}^{2}-2(a^{2}+r^{2})\,\partial_{v}\partial_{r}-2a\,\partial_{v}\partial_{\psi}-2a\,\partial_{r}\partial_{\psi}-\frac{1}{1-z^{2}}\,\partial_{\psi}^{2}-\Delta\,\partial_{r}^{2}\\ &+2(iasz-r)\,\partial_{v}+2\left[(s-1)(r-M)-\frac{2sr\Delta}{a^{2}+r^{2}}\right]\partial_{r}-\partial_{z}\!\left[(1-z^{2})\,\partial_{z}\right]\\ &-2s\left(\frac{2ar}{a^{2}+r^{2}}+\frac{iz}{1-z^{2}}\right)\partial_{\psi}+\left(\frac{s^{2}z^{2}}{1-z^{2}}-s+8Mr\,\frac{(s-2)a^{2}+2r^{2}}{(a^{2}+r^{2})^{2}}\right)\Bigg)\Psi^{(s)}=4\pi\Sigma\,T^{(s)}.\end{split}$ (2.12)
Here $\Psi^{(s)}$ denotes the perturbation field of spin weight $s$, and $T^{(s)}$ is the associated source term, which vanishes in vacuum. This form is obtained by projecting the Teukolsky operator given in Eq. (3.28) of Ref. [Cano:2023tmv] onto the HH tetrad. Equation (2.12) therefore applies to gravitational perturbations, corresponding to spin weights $s=\pm 2$. For $s=\pm 2$, the Weyl scalars $\psi_{0}$ and $\zeta^{4}\psi_{4}$ provide particular solutions of the TME with spin weights +2+2 and 2-2, respectively. We note that the expression of the source term $T^{(s)}$ is also tetrad dependent and therefore its explicit expression differs from that given in Ref. [Teukolsky:1972my]. However, in this work, we focus on vacuum perturbations and set the source term to zero.
A remarkable feature of Eq. (2.12) is that it admits separation of variables. Each spin-weighted field $\Psi^{(s)}$ can be expanded as
$\Psi^{(s)}(v,r,z,\psi)=\int d\omega\sum_{\ell=|s|}^{\infty}\sum_{m=-\ell}^{\ell}e^{-i\omega v+im\psi}\,c_{\ell m}^{(s)}(\omega)\,R_{\omega\ell m}^{(s)}(r)\,S_{\omega\ell m}^{(s)}(z)\,$, (2.13)
where the radial mode functions $R_{\omega\ell m}^{(s)}(r)$ and the angular mode functions $S_{\omega\ell m}^{(s)}(z)$ satisfy separate second-order ordinary differential equations,
$\Bigg[$ $\Delta\frac{d^{2}}{dr^{2}}-2\left((s-1)(r-M)-\frac{2sr\Delta}{a^{2}+r^{2}}+iK\right)\frac{d}{dr}-\mathscr{V}_{\rm rad}(r)-\lambda^{(s)}_{\omega\ell m}\Bigg]R_{\omega\ell m}^{(s)}(r)=0,$ (2.14a)
$\Bigg[$ $\frac{d}{dz}\left((1-z^{2})\frac{d}{dz}\right)+a^{2}\omega^{2}z^{2}-\frac{(m+sz)^{2}}{1-z^{2}}-2a\omega sz+s+A\Bigg]S_{\omega\ell m}^{(s)}(z)=0.$ (2.14b)
Here $\lambda^{(s)}_{\omega\ell m}$ is the separation constant, which must in general be determined numerically, and we define
$\begin{split}K&\equiv(r^{2}+a^{2})\omega-am,\qquad A\equiv\lambda^{(s)}_{\omega\ell m}+2am\omega-(a\omega)^{2},\\ K_{s}&\equiv K-(2s-1)am,\qquad\mathscr{V}_{\rm rad}(r)\equiv\frac{2irK_{s}}{a^{2}+r^{2}}+8Mr\frac{(s-2)a^{2}+2r^{2}}{(a^{2}+r^{2})^{2}}.\end{split}$ (2.15)
For $\omega\neq 0$, the angular equation defines the spin-weighted spheroidal harmonics.
2.1 Static modes
We restrict our analysis to slowly varying tidal fields, so that the TME (2.12) can be solved for static modes, i.e., with $\omega=0$. In this case, Eq. (2.14b) for the angular mode functions is satisfied by the spin-weighted spherical harmonics $_{s}Y_{\ell m}(z,\psi)$, and the separation constant reduces to [vandeMeent:2015lxa, Berens:2024czo]
$\lambda_{\ell m}^{(s)}\equiv\lambda_{0\ell m}^{(s)}=\ell(\ell+1)-s(s+1).$ (2.16)
The spin-raising operator $\eth_{s}$ and the spin-lowering operator $\bar{\eth}_{s}$ are defined as
$\eth_{s}$ $\equiv\frac{1}{\sqrt{1-z^{2}}}\left((1-z^{2})\,\partial_{z}-i\,\partial_{\psi}+sz\right),$ (2.17)
$\bar{\eth}_{s}$ $\equiv\frac{1}{\sqrt{1-z^{2}}}\left((1-z^{2})\,\partial_{z}+i\,\partial_{\psi}-sz\right).$ (2.18)
They allow to construct the spin-weighted spherical harmonics $_{s}Y_{\ell m}$ recursively from the scalar spherical harmonics $Y_{\ell m}\equiv{}_{0}Y_{\ell m}$, for $\ell\geq|s|$, according to
${}_{s+1}Y_{\ell m}$ $\equiv\left[(\ell-s)(\ell+s+1)\right]^{-1/2}\,\eth_{s}({}_{s}Y_{\ell m}),$ $\qquad\text{for $s\geq 0$},$ (2.19)
${}_{s-1}Y_{\ell m}$ $\equiv-\left[(\ell+s)(\ell-s+1)\right]^{-1/2}\,\bar{\eth}_{s}({}_{s}Y_{\ell m}),$ $\qquad\text{for $s\leq 0$}.$ (2.20)
Here
$Y_{\ell m}(z,\psi)\equiv\sqrt{\frac{(2\ell+1)(\ell-m)!}{4\pi(\ell+m)!}}\,P_{\ell}^{m}(z)\,e^{im\psi}=(-1)^{m}\,\bar{Y}_{\ell,-m}(z,\psi),$ (2.21)
where $P_{\ell}^{m}(z)$ are the associated Legendre polynomials of degree $\ell$ and order $m$. The spin-weighted spherical harmonics $_{s}Y_{\ell m}(z,\psi)$ form an orthonormal basis on the unit two-sphere $\mathbb{S}^{2}$, satisfying the orthogonality relation
$\oint_{\mathbb{S}^{2}}{}_{s}Y_{\ell m}\,{}_{s}\bar{Y}_{\ell^{\prime}m^{\prime}}\,d\Omega=\delta_{\ell\ell^{\prime}}\,\delta_{mm^{\prime}}.$ (2.22)
The radial equation (2.14a) for static modes, $R^{(s)}_{\ell m}\equiv R^{(s)}_{0\ell m}$, then reduces to
$x(x+1)\,\frac{d^{2}R^{(s)}_{\ell m}}{dx^{2}}+\left[2i\gamma m-(2x+1)(s-1)+\frac{4sx(x+1)(\delta+x)}{\gamma^{2}+(\delta+x)^{2}}\right]\frac{dR^{(s)}_{\ell m}}{dx}$
$\quad+\left[\frac{4(\delta+x)\bigl(i\gamma ms-4\delta+2\bigr)\bigl(\gamma^{2}+(\delta+x)^{2}\bigr)-\gamma^{2}(2\delta-1)(s-4)}{\bigl(\gamma^{2}+(\delta+x)^{2}\bigr)^{2}}-\lambda^{(s)}_{\ell m}\right]R^{(s)}_{\ell m}(x)=0\,.$ (2.23)
Here, we have introduced the dimensionless variables
$x\equiv\frac{r-r_{+}}{r_{+}-r_{-}},\qquad\gamma\equiv\frac{a}{r_{+}-r_{-}},\qquad\delta\equiv\frac{r_{+}}{r_{+}-r_{-}}.$ (2.24)
As already mentioned, we restrict our analysis to the gravitational case $s=\pm 2$.
3 Metric reconstruction
The reconstruction of the metric perturbation from solutions of the TME relies on the introduction of two tensorial differential operators, ${\mathscr{S}_{0}^{\dagger}}_{\mu\nu}$ and ${\mathscr{S}_{4}^{\dagger}}_{\mu\nu}$, which, when applied to appropriate scalar potentials, yield the corresponding metric perturbation in a radiation gauge. These scalar quantities, known as Hertz potentials, satisfy an adjoint version of the TME and serve as intermediaries between the Weyl scalars $\psi_{0}$, $\psi_{4}$, and the metric perturbation $h_{\mu\nu}$ [Chrzanowski:1975wv, Wald:1973wwa, Berens:2024czo, LeTiec:2020bos].
For a solution $\Psi_{\rm H}$ of the adjoint Teukolsky equation associated with the spin-s=2s=-2 field, the metric perturbation in the ingoing radiation gauge (IRG) is given by
$h_{\mu\nu}^{\rm IRG}=2\,\mathrm{Re}\!\left({\mathscr{S}_{0}^{\dagger}}_{\mu\nu}\,\Psi_{\rm H}\right)=2\,\mathrm{Re}\!\left({\mathscr{S}_{0}^{\dagger}}_{\mu\nu}\,\Psi^{(-2)}\right),$ (3.1)
which satisfies the gauge conditions $l^{\mu}h_{\mu\nu}^{\rm IRG}=0$ and $g^{\mu\nu}h_{\mu\nu}^{\rm IRG}=0$.
Similarly, for a Hertz potential $\Psi^{\prime}_{\rm H}$ associated with the spin-$s=+2$ field, the metric perturbation in the outgoing radiation gauge (ORG) reads
$h_{\mu\nu}^{\rm ORG}=2\,\mathrm{Re}\!\left({\mathscr{S}_{4}^{\dagger}}_{\mu\nu}\,\Psi^{\prime}_{\rm H}\right)=2\,\mathrm{Re}\!\left({\mathscr{S}_{4}^{\dagger}}_{\mu\nu}\,\zeta^{4}\Psi^{(+2)}\right),$ (3.2)
and satisfies $n^{\mu}h_{\mu\nu}^{\rm ORG}=0$ and $g^{\mu\nu}h_{\mu\nu}^{\rm ORG}=0$.
The Hertz potentials $\Psi_{\rm H}$ and $\Psi^{\prime}_{\rm H}$ are related to the Weyl scalars through fourth-order differential operators. In particular, the fields $\Psi^{(+2)}$ and $\Psi^{(-2)}$ appearing in Eqs. (3.2) and (3.1) are auxiliary Teukolsky fields of spin weight +2+2 and 2-2, respectively, and should not be confused with the Weyl scalars $\psi_{0}$ or $\zeta^{4}\psi_{4}$ [vandeMeent:2015lxa, LeTiec:2020bos, Berens:2024czo].
[…]
In the Newman–Penrose formalism, the connection associated with the null tetrad is encoded in twelve complex spin coefficients $(\kappa,\rho,\sigma,\tau,\kappa^{\prime},\rho^{\prime},\sigma^{\prime},\tau^{\prime},\epsilon,\beta,\epsilon^{\prime},\beta^{\prime})$. In the GHP formalism, these coefficients enter naturally through the derivative operators, which act on quantities of GHP weight $(p,q)$ according to
$\text{\thorn}=\mathscr{D}-p\epsilon-q\overline{\epsilon},\qquad\eth=\delta-p\beta+q\overline{\beta}^{\prime},\qquad\text{\thorn}^{\prime}=\mathscr{D}^{\prime}+p\overline{\epsilon}+q\overline{\epsilon}^{\prime},\qquad\eth^{\prime}=\delta^{\prime}+p\beta^{\prime}-q\overline{\beta},$ (3.3)
where the frame derivatives are defined as
$\mathscr{D}=l^{\mu}\nabla_{\mu},\qquad\mathscr{D}^{\prime}=n^{\mu}\nabla_{\mu},\qquad\delta=m^{\mu}\nabla_{\mu},\qquad\delta^{\prime}=\overline{m}^{\mu}\nabla_{\mu}.$ (3.4)
In this notation, the reconstruction operators ${\mathscr{S}_{0}^{\dagger}}_{\mu\nu}$ and ${\mathscr{S}_{4}^{\dagger}}_{\mu\nu}$ used in Eqs. (3.1) and (3.2) read [Berens:2024czo]
${\mathscr{S}_{0}^{\dagger}}_{\mu\nu}\Psi$ $=-\frac{1}{2}\,l_{\mu}l_{\nu}\bigl(\eth-\tau\bigr)\bigl(\eth+3\tau\bigr)\Psi-\frac{1}{2}\,m_{\mu}m_{\nu}\bigl(\text{\thorn}-\rho\bigr)\bigl(\text{\thorn}+3\rho\bigr)\Psi$
$\quad+\frac{1}{2}\,l_{(\mu}m_{\nu)}\Bigl[\bigl(\text{\thorn}-\rho+\overline{\rho}\bigr)\bigl(\eth+3\tau\bigr)+\bigl(\eth-\tau+\overline{\tau}^{\prime}\bigr)\bigl(\text{\thorn}+3\rho\bigr)\Bigr]\Psi,$ (3.5a)
${\mathscr{S}_{4}^{\dagger}}_{\mu\nu}\Psi$ $=-\frac{1}{2}\,n_{\mu}n_{\nu}\bigl(\eth^{\prime}-\tau^{\prime}\bigr)\bigl(\eth^{\prime}+3\tau^{\prime}\bigr)\Psi-\frac{1}{2}\,\overline{m}_{\mu}\overline{m}_{\nu}\bigl(\text{\thorn}^{\prime}-\rho^{\prime}\bigr)\bigl(\text{\thorn}^{\prime}+3\rho^{\prime}\bigr)\Psi$
$\quad+\frac{1}{2}\,n_{(\mu}\overline{m}_{\nu)}\Bigl[\bigl(\text{\thorn}^{\prime}-\rho^{\prime}+\overline{\rho}^{\prime}\bigr)\bigl(\eth^{\prime}+3\tau^{\prime}\bigr)+\bigl(\eth^{\prime}-\tau^{\prime}+\overline{\tau}\bigr)\bigl(\text{\thorn}^{\prime}+3\rho^{\prime}\bigr)\Bigr]\Psi.$ (3.5b)
In the HH tetrad (2.10), the relevant spin coefficients are
$\tau$ $=-\frac{ia\sqrt{1-z^{2}}}{\sqrt{2}\,\Sigma},\qquad\tau^{\prime}=-\frac{ia\sqrt{1-z^{2}}}{\sqrt{2}\,\zeta^{2}},\qquad\rho=\frac{\Delta}{2(r^{2}+a^{2})\,\zeta},\qquad\rho^{\prime}=\frac{r^{2}+a^{2}}{\zeta\,\Sigma},$ (3.6a)
$\beta$ $=\frac{z}{2\sqrt{2(1-z^{2})}\,\overline{\zeta}},\qquad\beta^{\prime}=\frac{rz+ia(z^{2}-2)}{2\sqrt{2(1-z^{2})}\,\zeta^{2}},\qquad\epsilon=\frac{M(r^{2}-a^{2})}{2(r^{2}+a^{2})^{2}},\qquad\epsilon^{\prime}=\frac{a(a+irz)}{\zeta\,\Sigma}.$ (3.6b)
The remaining spin coefficients $(\kappa,\kappa^{\prime},\sigma,\sigma^{\prime})$ vanish identically in the HH tetrad. In the following, we work with perturbations in the ORG, Eq. (3.2). This choice is standard for computing physically meaningful quantities, as in this gauge the metric perturbation is transverse at future null infinity $\mathscr{I}^{+}$ and naturally isolates the outgoing gravitational radiation [Chrzanowski:1975wv, vandeMeent:2015lxa].
3.1 Hertz potential
In the ORG, the relevant solution of the TME corresponds to a Hertz potential $\Psi^{\prime}_{\rm H}$ with spin weight $s=+2$.
In this work, we restrict our analysis to the lowest multipolar order of the external tidal field, namely the quadrupole order corresponding to $\ell=2$. The Hertz potential $\Psi^{\prime}_{\rm H}$ can therefore be decomposed as
$\Psi^{\prime}_{\rm H}=\sum_{m=-2}^{2}\Psi_{m}^{(+2)}.$ (3.7)
Each $m$-mode takes the general form
$\Psi_{m}^{(+2)}=c_{2m}(\tilde{v},a)\,\zeta^{4}\,R_{2m}(r)\,{}_{2}Y_{2m}(z,\psi),$ (3.8)
where $R_{2m}(r)$ are radial mode functions solving Eq. (2.14a) with $\omega=0$, $s=2$, and $\ell=2$, $_{2}Y_{2m}(z,\psi)$ are spin-weighted spherical harmonics, and $c_{2m}(\tilde{v},a)$ are complex coefficients encoding the external tidal field.
Here $\tilde{v}$ denotes a slowly varying time variable, in the sense that the tidal field is assumed to vary on timescales much longer than the characteristic dynamical timescale of the black hole. As a consequence, time derivatives are negligible compared to spatial derivatives [Binnington:2009bb, LeTiec:2020bos]. Accordingly, the Teukolsky equation is solved in the static ($\omega=0$) sector, and the coefficients $c_{2m}$ depend on $\tilde{v}$ only parametrically.
[…]
In Eq. (3.8), the radial mode function is given by
$R_{2m}(r)=\frac{\Delta^{2}}{4(r^{2}+a^{2})^{2}}\,\frac{\Gamma(3+2im\gamma)}{4!}\,\frac{\mathbf{F}(-4,1,-1+2im\gamma;-x)}{x^{2}(1+x)^{2}},\qquad|m|\leq 2,$ (3.9)
where $\mathbf{F}(a,b,c;z)$ denotes the regularized hypergeometric function, and $x$ and $\gamma$ are defined in Eq. (2.24). This expression coincides with the result obtained in the Kinnersley tetrad [LeTiec:2020bos], up to the expected rescaling by the square of the factor $\Delta/[2(r^{2}+a^{2})]$ relating the Kinnersley and HH tetrads, in agreement with Eq. (2.11).
The spin-weighted spherical harmonics $_{2}Y_{2m}(z,\psi)$ appearing in Eq. (3.8) follow from Eq. (2.19) with $s=2$ and $\ell=2$, and read
$_{2}Y_{2m}(z,\psi)=\frac{1}{2\sqrt{6}}\,\eth_{1}\eth_{0}\,Y_{2m}(z,\psi),\qquad|m|\leq 2,$ (3.10)
where the spin-raising operator $\eth_{s}$ is defined in Eq. (2.17), and $Y_{2m}(z,\psi)$ are the standard spherical harmonics given in Eq. (2.21). Inserting this expression together with Eq. (3.9) into Eq. (3.8) yields
$\Psi^{(+2)}_{m}=c_{2m}(\tilde{v})\,(r-iaz)^{4}\,\frac{(r_{+}-r_{-})^{4}}{4(r^{2}+a^{2})^{2}}\,\frac{\Gamma(3+2im\gamma)}{4!}\,\mathbf{F}(-4,1,-1+2im\gamma;-x)\,{}_{2}Y_{2m}(z,\psi).$ (3.11)
The asymptotic behavior of the reconstructed metric must correspond to that of a tidally deformed black hole. This requires identifying the Weyl tensor in the intermediate region $r_{+}\ll r\ll\mathscr{R}$ with the Weyl tensor of the external background spacetime responsible for the tidal deformation. Here $\mathscr{R}$ denotes the curvature scale of the external spacetime.
From the perspective of the external spacetime, the black hole can be treated as a point particle moving along a worldline $\mathscr{L}$, which is endowed with an orthonormal tetrad $(u^{\alpha},e^{\alpha}_{a})$. Here $u^{\alpha}$ is the black-hole four-velocity and $e^{\alpha}_{a}$ is a spatial triad parallel transported along $\mathscr{L}$. This tetrad defines the local asymptotic rest frame of the black hole within the external geometry. Projecting the external Weyl tensor onto this frame yields the components [Zhang:1986cpa, Poisson:2009qj, Binnington:2009bb, Chatziioannou:2012gq, Poisson:2014gka]
$C_{0b0d}\equiv C_{\alpha\beta\gamma\delta}u^{\alpha}e^{\beta}_{b}u^{\gamma}e^{\delta}_{d},\qquad C_{0bcd}\equiv C_{\alpha\beta\gamma\delta}u^{\alpha}e^{\beta}_{b}e^{\gamma}_{c}e^{\delta}_{d}.$ (3.12)
These frame components can in turn be recast in terms of two symmetric, trace-free tensors $\mathscr{E}_{ab}$ and $\mathscr{B}_{ab}$, representing the electric-type and magnetic-type quadrupole moments of the tidal field, respectively [Poisson:2009qj]. They are defined as
$\mathscr{E}_{ab}\equiv C_{0a0b},\qquad\mathscr{B}_{ab}\equiv\frac{1}{2}\,\epsilon_{a}{}^{cd}\,C_{cdb0},$ (3.13)
where $\epsilon_{abc}$ denotes the three-dimensional Levi–Civita tensor. Together, these tensors encode all ten independent components of the Weyl tensor and therefore provide a complete description of the tidal environment acting on the black hole.
[…]
As an alternative to the standard asymptotic-matching approach of Refs. [Poisson:2009qj, Binnington:2009bb, Poisson:2004cw, Chatziioannou:2012gq], we determine the coefficients $c_{2m}$ by requiring that the metric reconstructed from $\Psi^{\prime}_{\rm H}$ reproduces the known quadrupolar tidal perturbation computed in Ref. [Binnington:2009bb] in the Schwarzschild limit $a=0$. The two procedures are equivalent and consistent with each other once the relation between the Weyl scalar $\psi_{0}$ and the Hertz potential $\Psi^{\prime}_{\rm H}$ is taken into account (see, e.g., Eq. (1.23a) of Ref. [Berens:2024czo]). The details of our procedure, along with the final expressions for $c_{2m}$, are presented at the end of Sec. 4.
4 Perturbation to the Kerr metric: results
In this section, we present the explicit form of the quadrupolar perturbation to the Kerr metric reconstructed in the ORG. The perturbation is obtained by applying the reconstruction operator ${\mathscr{S}_{4}^{\dagger}}_{\mu\nu}$, defined in Eq. (3.5), to the Hertz potential with spin weight $s=+2$. For each value of the azimuthal number $m$, we define the complex quantity
$\hat{h}_{\mu\nu}^{(m)\,\rm ORG}\equiv{\mathscr{S}_{4}^{\dagger}}_{\mu\nu}\,\Psi^{(+2)}_{m}$, (4.1)
where $\Psi^{(+2)}_{m}$ is defined in Eq. (3.11).
The full complex metric perturbation is then obtained by summing over the allowed values of $m$,
$\hat{h}_{\mu\nu}^{\rm ORG}=\sum_{m=-2}^{2}\hat{h}_{\mu\nu}^{(m)\,\rm ORG}.$ (4.2)
The final step, as indicated in Eq. (3.2), is to take the real part of this expression, yielding $h_{\mu\nu}^{\rm ORG}\equiv 2\,\mathrm{Re}\!\left[\hat{h}_{\mu\nu}^{\rm ORG}\right]$. The explicit expressions for each $m$-mode are given below. To simplify the notation, we henceforth omit the superscript “ORG”. We present the full expressions explicitly in order to facilitate direct reuse in analytic and numerical applications.
4.1 $m=0$ mode
We begin with the axisymmetric contribution $m=0$, which corresponds to the purely polar part of the quadrupolar deformation. The corresponding components $\hat{h}_{\mu\nu}^{(0)}$ are listed below.
$\hat{h}_{vv}^{(0)}$ $=\frac{\mathscr{N}_{2\,0}}{\zeta\overline{\zeta}^{3}}\Bigg[a^{6}z^{4}(z^{2}-3)+2a^{5}iz(z^{2}-1)\left(M(3z^{2}+1)-3rz^{2}\right)$
$+a^{4}\left(2M^{2}z^{2}(z^{2}-1)^{2}-2Mr(3z^{6}-9z^{4}+3z^{2}-1)+3r^{2}z^{4}(z^{2}-3)\right)$
$-2a^{3}irz(z^{2}-1)\left(M^{2}(6z^{2}+2)-3Mr(3z^{2}+1)+3r^{2}(z^{2}+1)\right)$
$-a^{2}r^{2}\left(2M^{2}(3z^{4}+1)-4Mr(z^{4}+4z^{2}-1)+3r^{2}(3z^{2}-1)\right)$
$-2air^{3}z(z^{2}-1)(4M^{2}-8Mr+3r^{2})-r^{4}(3z^{2}-1)(2M-r)^{2}\Bigg]\,$, (4.3a)
$\hat{h}_{vz}^{(0)}$ $=-\frac{i~\mathscr{N}_{2\,0}}{\overline{\zeta}^{2}}\Bigg[a^{5}z^{2}(z^{2}-1)+2ia^{4}z\left(2Mz^{2}-r(2z^{2}-1)\right)+a^{3}(z^{2}-1)\left(2M^{2}z^{2}-r(2M-r)(3z^{2}-1)\right)$
$-4ia^{2}rz\left(2M^{2}z^{2}-Mr(3z^{2}-1)+r^{2}z^{2}\right)-ar^{2}(z^{2}-1)\left(2M^{2}-r^{2}\right)+2ir^{4}z(2M-r)\Bigg]\,$, (4.3b)
$\hat{h}_{v\psi}^{(0)}$ $=2\frac{\mathscr{N}_{2\,0}}{\overline{\zeta}^{3}}(1-z^{2})\Bigg[ia^{6}z^{3}-a^{5}\left(z^{4}(r-M)+z^{2}(2M-3r)+M\right)$
$+ia^{4}z\left(M^{2}(z^{2}-1)-Mr(6z^{2}-4)+r^{2}(4z^{2}-3)\right)$
$-a^{3}r\left(M^{2}(2z^{4}-5z^{2}-1)-Mr(3z^{4}-12z^{2}+1)+r^{2}(z^{4}-6z^{2}+1)\right)$
$ia^{2}r^{2}z\left(5M^{2}(z^{2}-1)-8Mr(z^{2}-1)+r^{2}(3z^{2}-4)\right)$
$+a\left(r^{3}(M-r)(z^{2}(5M-3r)-M+r)\right)+ir^{5}z(2M-r)\Bigg]\,$, (4.3c)
$\hat{h}_{vr}^{(0)}$~ $=~\hat{h}_{rr}^{(0)}~=~\hat{h}_{rz}^{(0)}~=~\hat{h}_{r\psi}^{(0)}~=~0\,$, (4.3d)
$\hat{h}_{zz}^{(0)}$ $=\mathscr{N}_{2\,0}\Bigg[a^{4}z^{2}+2ia^{3}Mz+a^{2}\left(2M^{2}z^{2}-r(2M-r)(3z^{2}+1)\right)-2iar^{2}(3M-2r)z$
$+r^{2}\left(2M^{2}-r^{2}\right)\Bigg]\,$, (4.3e)
$\hat{h}_{z\psi}^{(0)}$ $=-\frac{\mathscr{N}_{2\,0}}{\overline{\zeta}^{2}}(1-z^{2})\Bigg[ia^{6}z^{2}+2a^{5}\left(Mz^{3}-rz(z^{2}-1)\right)+ia^{4}\left(2M^{2}z^{2}-2Mr(3z^{2}-1)+r^{2}(4z^{2}-1)\right)$
$-a^{3}\left(4M^{2}rz^{3}-2r^{2}(3M-r)z(z^{2}-1)\right)+ia^{2}r^{2}\left(2M^{2}(z^{2}-1)-2Mr(3z^{2}-1)+3r^{2}z^{2}\right)$
$+2aMr^{3}z(2M-r)-ir^{4}(2M^{2}-r^{2})\Bigg]\,$, (4.3f)
$\hat{h}_{\psi\psi}^{(0)}$ $=-\frac{\mathscr{N}_{2\,0}}{\overline{\zeta}^{3}}(1-z^{2})^{2}\Bigg[ia^{7}z^{3}-a^{6}\left(2M(2z^{2}+1)-3rz^{2}\right)-ia^{5}z\left(2M^{2}+r(2M-r)(2z^{2}-3)\right)$
$+a^{4}\left(2M^{2}r(4z^{2}+1)-12Mr^{2}z^{2}+r^{3}(6z^{2}-1)\right)+ia^{3}\left(4Mr^{2}(M-r)z(z^{2}-2)+r^{4}z(z^{2}-6)\right)$
$+a^{2}r^{3}\left(4M^{2}r^{3}z^{2}-2Mr^{4}(4z^{2}-1)+r^{5}(3z^{2}-2)\right)+iar^{4}(6M^{2}+2Mr-3r^{2})z+r^{5}(2M^{2}-r^{2})\Bigg]\,$, (4.3g)
where we defined the complex coefficient
$\mathscr{N}_{2\,0}\equiv\frac{c_{2\,0}}{16}\sqrt{\frac{15}{2\pi}}\,.$ (4.4)
4.2 $m=\pm 1$ modes
These modes are obtained analogously from Eq. (4.1) with $m=\pm 1$, and their explicit expressions are reported below.
$\hat{h}_{vv}^{(\pm 1)}$ $=-\frac{\mathscr{N}_{2\,\pm 1}}{{\overline{\zeta}^{3}}}\sqrt{1-\mathbf{z}^{2}}\Bigg[9\mathbf{i}a^{5}\mathbf{z}^{3}+3a^{4}\left(M(3\mathbf{z}^{4}-3\mathbf{z}^{3}-5\mathbf{z}^{2}+\mathbf{z}-2)-3r\mathbf{z}^{2}(\mathbf{z}^{2}-\mathbf{z}-3)\right)$
$+3\mathbf{i}a^{3}\Big(2M^{2}(\mathbf{z}^{2}+\mathbf{z}^{3}-2\mathbf{z}^{2}-1)+3r^{2}\mathbf{z}(\mathbf{z}^{3}+3\mathbf{z}^{2}-3\mathbf{z}-3)\Big)$
$+a^{2}\Big(4M^{3}(\mathbf{z}^{2}+\mathbf{z}+1)+6M^{2}r(7\mathbf{z}^{3}+\mathbf{z}^{2}-4\mathbf{z}-1)-18Mr^{2}(4\mathbf{z}^{3}+3\mathbf{z}^{2}-3\mathbf{z}-1)$
$+9r^{3}(3\mathbf{z}^{3}+3\mathbf{z}^{2}-3\mathbf{z}-1)\Big)$
$-3\mathbf{i}ar\Big(4M^{3}\mathbf{z}+4M^{2}r(5\mathbf{z}^{2}+\mathbf{z}-1)-2Mr^{2}(14\mathbf{z}^{2}+5\mathbf{z}-4)+3r^{3}(3\mathbf{z}^{2}+\mathbf{z}-1)\Big)$
$-9r^{3}(2M-r)^{2}\mathbf{z}\Bigg]\,$, (4.5a)
$\hat{h}_{vz}^{(\pm 1)}$ $=\pm\frac{\mathscr{N}_{2\,\pm 1}}{{\overline{\zeta}^{2}}}\sqrt{\frac{1-\mathbf{z}}{1+\mathbf{z}}}\Bigg[3\mathbf{i}a^{5}\mathbf{z}^{2}(2\mathbf{z}+1)+3a^{4}\mathbf{z}\left(3M\mathbf{z}(\mathbf{z}^{2}-1)-r(3\mathbf{z}^{3}-4\mathbf{z}-2)\right)$
$+\mathbf{i}a^{3}\left(2M^{2}\mathbf{z}^{2}(3\mathbf{z}^{2}+4\mathbf{z}-1)-6Mr(3\mathbf{z}^{4}+6\mathbf{z}^{3}-2\mathbf{z}-1)+3r^{2}(3\mathbf{z}^{4}+6\mathbf{z}^{3}-2\mathbf{z}-1)\right)$
$+a^{2}r\left(2M^{2}(12\mathbf{z}^{3}+8\mathbf{z}^{2}-2\mathbf{z}-3)-9Mr(4\mathbf{z}^{3}+3\mathbf{z}^{2}-1)+6r^{2}\mathbf{z}^{2}(2\mathbf{z}+1)\right)$
$-\mathbf{i}ar\left(4M^{3}(2\mathbf{z}+1)+2M^{2}r(3\mathbf{z}^{2}+4\mathbf{z}-1)-3r^{3}\mathbf{z}(\mathbf{z}+2)\right)$
$-3r^{4}(2M-r)(2\mathbf{z}+1)\Bigg]\,$, (4.5b)
$\hat{h}_{v\psi}^{(\pm 1)}$ $=\frac{\mathscr{N}_{2\,\pm 1}}{{\overline{\zeta}^{3}}}(1-\mathbf{\mathbf{z}})\sqrt{1-\mathbf{\mathbf{z}}^{2}}\Bigg[3\mathbf{i}a^{6}\mathbf{z}^{3}(\mathbf{z}+2)-3a^{5}(\mathbf{z}+1)\left(M(5\mathbf{z}^{2}-\mathbf{z}+2)-6r\mathbf{z}^{2}\right)$
$+\mathbf{i}a^{4}\left(2M^{2}(2\mathbf{z}^{4}-2\mathbf{z}^{3}-6\mathbf{z}^{2}-3\mathbf{z}-3)-3Mr(6\mathbf{z}^{4}-17\mathbf{z}^{2}-10\mathbf{z}-3)+9r^{2}\mathbf{z}(\mathbf{z}^{3}-4\mathbf{z}-2)\right)$
$+a^{3}\Big(4M^{3}(\mathbf{z}^{3}+2\mathbf{z}^{2}+2\mathbf{z}+1)+6M^{2}r(2\mathbf{z}^{4}+4\mathbf{z}^{3}-2\mathbf{z}^{2}-4\mathbf{z}-1)$
$-3Mr^{2}(6\mathbf{z}^{4}+21\mathbf{z}^{3}-17\mathbf{z}-4)+6r^{3}(\mathbf{z}^{4}+5\mathbf{z}^{3}-5\mathbf{z}-1)\Big)$
$-\mathbf{i}a^{2}r\Big(4M^{3}\mathbf{z}(\mathbf{z}+2)+6M^{2}r(5\mathbf{z}^{3}+8\mathbf{z}^{2}+\mathbf{z}-1)-3Mr^{2}(16\mathbf{z}^{3}+29\mathbf{z}^{2}+2\mathbf{z}-5)$
$+9r^{3}(2\mathbf{z}^{3}+4\mathbf{z}^{2}-1)\Big)$
$+2ar^{2}\Big(2M^{3}(2\mathbf{z}+1)-M^{2}r(15\mathbf{z}^{2}+14\mathbf{z}+1)+3Mr^{2}\mathbf{z}(8\mathbf{z}+7)-9r^{3}\mathbf{z}(\mathbf{z}+1)\Big)$
$-3\mathbf{i}r^{5}(2M-r)(2\mathbf{z}+1)\Bigg]\,$, (4.5c)
$\hat{h}_{vr}^{(\pm 1)}$~ $=~\hat{h}_{rr}^{(\pm 1)}~=~\hat{h}_{rz}^{(\pm 1)}~=~\hat{h}_{r\psi}^{(\pm 1)}=0\,$, (4.5d)
$\hat{h}_{zz}^{(\pm 1)}$ $=-\mathbf{i}\mathscr{N}_{2\,\pm 1}\zeta\sqrt{\frac{1-\mathbf{z}}{1+\mathbf{z}}}\Bigg[3a^{3}-9\mathbf{i}a^{2}\mathbf{z}(M-r)+a\left(M^{2}(6\mathbf{z}+4)-9M(2r\mathbf{z}+r)+9r^{2}\mathbf{z}\right)$
$-6\mathbf{i}M^{2}r+3\mathbf{i}r^{3}\Bigg]\,$, (4.5e)
$\hat{h}_{z\psi}^{(\pm 1)}$ $=\mp\frac{\mathscr{N}_{2\,\pm 1}}{{\overline{\zeta}^{3}}}(1-\mathbf{\mathbf{z}})\sqrt{1-\mathbf{\mathbf{z}}^{2}}\Bigg[3\mathbf{i}a^{6}\mathbf{z}^{2}(\mathbf{z}+1)-3a^{5}\mathbf{z}(3M\mathbf{z}-r(3\mathbf{z}+2))$
$+\mathbf{i}a^{4}\left(2M^{2}\mathbf{z}^{2}(2\mathbf{z}-1)-6Mr(3\mathbf{z}^{3}-2\mathbf{z}-1)+3r^{2}(3\mathbf{z}^{3}-3\mathbf{z}-1)\right)$
$+a^{3}r\left(2M^{2}(6\mathbf{z}^{3}+6\mathbf{z}^{2}-2\mathbf{z}-3)-9Mr(2\mathbf{z}^{3}+3\mathbf{z}^{2}-1)+3r^{2}(2\mathbf{z}^{3}+5\mathbf{z}^{2}-1)\right)$
$-\mathbf{i}a^{2}r\left(M^{3}(8\mathbf{z}+4)+2M^{2}r(3\mathbf{z}^{2}+6\mathbf{z}-1)-18Mr^{2}\mathbf{z}(\mathbf{z}+1)+3r^{3}\mathbf{z}(3\mathbf{z}+1)\right)$
$-ar^{3}\left(4M^{2}(3\mathbf{z}+1)-3Mr(2\mathbf{z}+1)-3r^{2}\right)+3\mathbf{i}r^{4}\left(2M^{2}-r^{2}\right)\Bigg]\,$, (4.5f)
$\hat{h}_{\psi\psi}^{(\pm 1)}$ $=\frac{\mathbf{i}\mathscr{N}_{2\,\pm 1}}{\zeta\overline{\zeta}^{2}}\sqrt{1-\mathbf{z}^{2}}\Bigg[(1+\mathbf{z})(1-\mathbf{z})^{2}(a^{2}+r^{2})^{2}\Bigg(3a^{3}-9\mathbf{i}a^{2}\mathbf{z}(M-r)$
$+a\left(M^{2}(6\mathbf{z}+4)-9M(2r\mathbf{z}+r)+9r^{2}\mathbf{z}\right)-6\mathbf{i}M^{2}r+3\mathbf{i}r^{3}\Bigg)$
$+a^{2}(1-\mathbf{z}^{2})^{2}\Bigg(3a^{4}+6\mathbf{i}a^{3}(M-r)-2a^{2}M^{2}-2\mathbf{i}a(M-r)(2M^{2}+6Mr-3r^{2})$
$-3r^{2}(2M-r)^{2}\Bigg)\left(a(\mathbf{z}^{2}+\mathbf{z}+1)+3\mathbf{i}r\mathbf{z}\right)\Bigg]$
$+2a\frac{\mathscr{N}_{2,\pm 1}}{\zeta\overline{\zeta}^{3}}\sqrt{1-\mathbf{z}^{2}}(1-\mathbf{z})^{2}(1+\mathbf{z})(a^{2}+r^{2})\Bigg[3a^{5}\mathbf{z}(\mathbf{z}^{2}+\mathbf{z}+1)-3\mathbf{i}a^{4}(2r\mathbf{z}+r)$
$+a^{3}\left(4M^{2}\mathbf{z}(\mathbf{z}^{2}+\mathbf{z}+1)-3r(2M-r)(3\mathbf{z}^{3}+3\mathbf{z}^{2}+\mathbf{z}-1)\right)$
$-2\mathbf{i}a^{2}r\left(M^{2}(6\mathbf{z}^{3}+6\mathbf{z}^{2}+4\mathbf{z}-1)-9Mr\mathbf{z}(\mathbf{z}^{2}+\mathbf{z}+1)+3r^{2}\mathbf{z}(\mathbf{z}^{2}+\mathbf{z}+1)\right)$
$-ar(4M^{3}+4M^{2}r-3r^{3})(2\mathbf{z}+1)+3\mathbf{i}r^{4}(2M-r)(2\mathbf{z}+1)\Bigg]\,$, (4.5g)
where we defined the complex function of the angle $\psi$
$\mathscr{N}_{2\,\pm 1}\equiv\frac{c_{2\,\pm 1}}{48}\sqrt{\frac{5}{\pi}}e^{\mathbf{i}\psi}$ (4.6)
with
$\mathbf{z}\equiv\begin{cases}z&\mbox{if }m=+1\\ -z&\mbox{if }m=-1\end{cases}\,,\quad\mbox{and}\quad\mathbf{i}\equiv\begin{cases}i&\mbox{if }m=+1\\ -i&\mbox{if }m=-1\end{cases}\,.$ (4.7)
4.3 $m=\pm 2$ modes
Finally, the $m=\pm 2$ components, which carry the highest azimuthal dependence, are given explicitly below.
$\hat{h}_{vv}^{(\pm 2)}$ $=\frac{\mathscr{N}_{2\,\pm 2}}{\overline{\zeta}^{3}}(\mathbf{z}^{2}-1)\Bigg[9\mathbf{i}a^{5}\mathbf{z}^{3}-3a^{4}\left(M(6\mathbf{z}^{3}+6\mathbf{z}^{2}-2\mathbf{z}+2)-3r\mathbf{z}^{2}(2\mathbf{z}+3)\right)$
$-3\mathbf{i}a^{3}\left(2M^{2}(\mathbf{z}+1)(\mathbf{z}^{2}+\mathbf{z}-4)-6Mr(\mathbf{z}^{3}+5\mathbf{z}^{2}+\mathbf{z}+1)+3r^{2}\mathbf{z}(\mathbf{z}^{2}+6\mathbf{z}+3)\right)$
$-a^{2}\left(8M^{3}(\mathbf{z}+2)+6M^{2}r(\mathbf{z}+1)(7\mathbf{z}-5)-36Mr^{2}(7\mathbf{z}^{2}+2\mathbf{z}-5)+9r^{3}(3\mathbf{z}^{2}+6\mathbf{z}+1)\right)$
$+3\mathbf{i}ar\left(8M^{3}+4M^{2}r(5\mathbf{z}+2)-4Mr^{2}(7\mathbf{z}+5)+3r^{3}(3\mathbf{z}+2)\right)$
$+9r^{3}(2M-r)^{2}\Bigg]\,$, (4.8a)
$\hat{h}_{vz}^{(\pm 2)}$ $=\pm\mathbf{i}\frac{\mathscr{N}_{2\,\pm 2}}{\overline{\zeta}^{2}}(1-\mathbf{z})\Bigg[3a^{5}\mathbf{z}^{2}(3\mathbf{z}+1)-6\mathbf{i}a^{4}\mathbf{z}(-3M\mathbf{z}(\mathbf{z}+1)+3r\mathbf{z}(\mathbf{z}+1)+r)$
$-a^{3}\left(2M^{2}\mathbf{z}^{2}(3\mathbf{z}+5)-3(3\mathbf{z}^{3}+9\mathbf{z}^{2}+3\mathbf{z}+1)(2Mr-r^{2})\right)$
$+2\mathbf{i}a^{2}r\left(4M^{2}(3\mathbf{z}^{2}+\mathbf{z}-3)-9Mr(2\mathbf{z}^{2}+\mathbf{z}+1)+6r^{2}\mathbf{z}^{2}\right)$
$+ar\left(16M^{3}+2M^{2}r(3\mathbf{z}+5)-3r^{3}(\mathbf{z}+3)\right)-6\mathbf{i}r^{4}(2M-r)\Bigg]\,$, (4.8b)
$\hat{h}_{v\psi}^{(\pm 2)}$ $=2\frac{\mathscr{N}_{2\,2\pm 2}}{\overline{\zeta}^{3}}(1-\mathbf{z})^{2}(1+\mathbf{z})\Bigg[3\mathbf{i}a^{6}z^{3}-3a^{5}(z+1)\left(M(3z-1)z+M-3rz^{2}\right)$
$+\mathbf{i}a^{4}\left(M^{2}(-4z^{3}+3z^{2}+21z+12)+3Mr(6z^{3}+12z^{2}+5z+3)-9r^{2}z(z^{2}+3z+1)\right)$
$-a^{3}\Big(4M^{3}(z^{2}+3z+2)+3M^{2}r(2z^{3}+6z^{2}+z-5)$
$-3Mr^{2}(3z^{3}+18z^{2}+18z+5)+3r^{3}(z^{3}+9z^{2}+9z+1)\Big)$
$+\mathbf{i}a^{2}r\left(4M^{3}(z+3)+3M^{2}r(5z^{2}+11z+8)-3Mr^{2}(8z^{2}+21z+7)+9r^{3}(z^{2}+3z+1)\right)$
$-ar^{2}\left(8M^{3}-M^{2}r(15z+13)+6Mr^{2}(4z+3)-9r^{3}(z+1)\right)+3\mathbf{i}r^{5}(2M-r)\Bigg]\,$, (4.8c)
$\hat{h}_{vr}^{(\pm 2)}$~ $=~\hat{h}_{rr}^{(\pm 2)}~=~\hat{h}_{rz}^{(\pm 2)}~=~\hat{h}_{r\psi}^{(\pm 2)}~=~0\,$, (4.8d)
$\hat{h}_{zz}^{(\pm 2)}$ $=-\mathbf{i}\mathscr{N}_{2\,\pm 2}\zeta\left(\frac{1-\mathbf{z}}{1+\mathbf{z}}\right)\Bigg[a^{3}(9\mathbf{z}+6)+9\mathbf{i}a^{2}\left(2M(\mathbf{z}+1)-r(2\mathbf{z}+1)\right)$
$-a\left(2M^{2}(3\mathbf{z}+4)-18Mr(\mathbf{z}+1)+9r^{2}\mathbf{z}\right)+3\mathbf{i}r(2M^{2}-r^{2})\Bigg]\,$, (4.8e)
$\hat{h}_{z\psi}^{(\pm 2)}$ $=\mp\mathbf{i}\frac{\mathscr{N}_{2\,\pm 2}}{\overline{\zeta}^{2}}(1-\mathbf{z}^{2})\Bigg[3a^{6}\mathbf{z}^{2}(2\mathbf{z}+1)-6\mathbf{i}a^{5}\mathbf{z}(-3M\mathbf{z}(\mathbf{z}+1)+3r\mathbf{z}(\mathbf{z}+1)+r)$
$-a^{4}\left(+2M^{2}\mathbf{z}^{2}(4\mathbf{z}+5)-6Mr(6\mathbf{z}^{3}+9\mathbf{z}^{2}+4\mathbf{z}+1)+3r^{2}(6\mathbf{z}^{3}+12\mathbf{z}^{2}+6\mathbf{z}+1)\right)$
$+2\mathbf{i}a^{3}r\left(2M^{2}(3\mathbf{z}^{3}+6\mathbf{z}^{2}-4\mathbf{z}-6)-9Mr(\mathbf{z}+1)^{3}+3r^{2}(\mathbf{z}^{3}+5\mathbf{z}^{2}+3\mathbf{z}+1)\right)$
$+a^{2}r\left(16M^{3}(\mathbf{z}+1)+2M^{2}r(3\mathbf{z}^{2}+12\mathbf{z}+5)-18Mr^{2}(\mathbf{z}+1)^{2}+3r^{3}\mathbf{z}(3\mathbf{z}+2)\right)$
$-2\mathbf{i}ar^{3}\left(2M^{2}(3\mathbf{z}+2)-3Mr(\mathbf{z}+1)-3r^{2}\right)-6M^{2}r^{4}+3r^{6}\Bigg]\,$, (4.8f)
$\hat{h}_{\psi\psi}^{(\pm 2)}$ $=\frac{\mathscr{N}_{2\,\pm 2}}{\zeta\overline{\zeta}^{2}}(1-\mathbf{z}^{2})\Bigg[\mathbf{i}a^{2}(\mathbf{z}^{2}-1)^{2}(a(\mathbf{z}+2)+3\mathbf{i}r)\Big(3a^{4}+12\mathbf{i}a^{3}(M-r)$
$+2a^{2}(4M^{2}+18Mr-9r^{2})+4\mathbf{i}a(M-r)(2M^{2}+6Mr-3r^{2})+3r^{2}(2M-r)^{2}\Big)$
$+(1-\mathbf{z}^{2})^{2}(a^{2}+r^{2})^{2}\Bigg(3\mathbf{i}a^{3}(3\mathbf{z}+2)-9a^{2}\left(2M(\mathbf{z}+1)-r(2\mathbf{z}+1)\right)$
$-\mathbf{i}a\left(M^{2}(6\mathbf{z}+8)-18Mr(\mathbf{z}+1)+9r^{2}\mathbf{z}\right)-6M^{2}r+3r^{3}\Bigg)\Bigg]$
$+4\frac{\mathscr{N}_{2,\pm 2}}{\zeta\overline{\zeta}^{3}}(1-\mathbf{z})(1-\mathbf{z}^{2})^{2}(a^{2}+r^{2})a\Bigg[3a^{5}\mathbf{z}(\mathbf{z}+1)+3\mathbf{i}a^{4}\left(3M\mathbf{z}(\mathbf{z}+1)-3r\mathbf{z}(\mathbf{z}+1)+r\right)$
$-a^{3}\left(4M^{2}\mathbf{z}(\mathbf{z}+1)-6Mr(3\mathbf{z}(\mathbf{z}+1)+2)+3r^{2}(3\mathbf{z}^{2}+3\mathbf{z}+2)\right)$
$+\mathbf{i}a^{2}r\left(M^{2}(6\mathbf{z}^{2}+6\mathbf{z}-8)-9Mr(\mathbf{z}^{2}+\mathbf{z}+2)+3r^{2}\mathbf{z}(\mathbf{z}+1)\right)$
$+2ar\left(4M^{3}+4M^{2}r-3r^{3}\right)-3\mathbf{i}r^{4}(2M-r)\Bigg]\,$, (4.8g)
where, analogously to the case $m\pm 1$, we defined
$\mathscr{N}_{2\,\pm 2}\equiv\frac{c_{2\,\pm 2}}{96}\sqrt{\frac{5}{\pi}}e^{2\mathbf{i}\psi}$ (4.9)
with
$\mathbf{z}\equiv\begin{cases}z&\mbox{if }m=+2\\ -z&\mbox{if }m=-2\end{cases}\,,\quad\mbox{and}\quad\mathbf{i}\equiv\begin{cases}i&\mbox{if }m=+2\\ -i&\mbox{if }m=-2\end{cases}\,.$ (4.10)
The complete perturbation $h_{\mu\nu}^{\rm ORG}$, obtained by summing the five contributions above and taking the real part of the result, satisfies the ORG conditions. One can further verify that, in the zero-spin limit $a=0$, it reduces to the known quadrupolar tidal perturbation of the Schwarzschild metric derived in Ref. [Binnington:2009bb]. Moreover, by expanding the reconstructed metric to first order in the spin parameter $a$, the associated Weyl scalars $\psi_{0}$ and $\psi_{4}$ agree with those obtained in Refs. [Poisson:2014gka, Landry:2015zfa].
Finally, we have explicitly verified that each mode of the metric perturbation $\hat{h}_{\mu\nu}^{(m)\,\rm ORG}$ satisfies the linearized Einstein equations. For completeness, the full expressions for the metric perturbation components are provided in electronic form as a Mathematica Notebook in the supplementary material to this paper.
By performing a suitable gauge transformation from the ORG to the IRG and rotating from the HH to the Kinnersley tetrad, the tidally perturbed Kerr metric constructed here can be related to those obtained in Refs. [LeTiec:2020bos, Yunes:2005ve]. We also note that explicit reconstruction formulas in both the IRG and ORG within the Kinnersley tetrad were recently presented in Ref. [Berens:2024czo]. [^2] See Ref. [Berens:2025kkm] for an extension of this formalism to Kerr black holes embedded in a spacetime with a nonzero cosmological constant. Those expressions, however, were neither specialized to tidal perturbations nor given in the fully explicit analytic form derived here, which facilitates direct use in astrophysical applications (see Secs. 5 and 6).
4.4 Fixing the coefficients $c_{2m}$
We now determine the coefficients $c_{2m}$ entering the metric through the normalization factors $\mathscr{N}_{2\,0}$, $\mathscr{N}_{2\,\pm 1}$, and $\mathscr{N}_{2\,\pm 2}$ in Eqs. (4.4), (4.6), and (4.9).
As anticipated in Sec. 3.1, rather than following the standard asymptotic-matching procedure employed in Refs. [Poisson:2009qj, Binnington:2009bb, Poisson:2004cw, Chatziioannou:2012gq], we take the zero-spin limit ($a=0$) of the metric perturbation (3.2) and require that it reproduces the quadrupolar order of the tidally deformed metric derived in Ref. [Binnington:2009bb] for a Schwarzschild black hole. This fixes the coefficients $c_{2m}$ in terms of the external tidal moments.
Introducing the complex combination of the tidal tensors $\mathscr{C}_{ab}\equiv\mathscr{E}_{ab}-i\mathscr{B}_{ab}$, the coefficients read
$c_{2\,0}$ $=-4\sqrt{\frac{2\pi}{15}}\,\bigl(\mathscr{C}_{11}+\mathscr{C}_{22}\bigr),$ (4.11)
$c_{2\,\pm 1}$ $=\mp\frac{8}{3}\sqrt{\frac{\pi}{5}}\,\bigl(\mathscr{C}_{13}\mp i\,\mathscr{C}_{23}\bigr),$
$c_{2\,\pm 2}$ $=\frac{4}{3}\sqrt{\frac{\pi}{5}}\,\bigl(\mathscr{C}_{11}-\mathscr{C}_{22}\mp 2i\,\mathscr{C}_{12}\bigr).$
As anticipated in Sec. 3, once the relation between the Weyl scalar $\psi_{0}$ and the Hertz potential $\Psi^{\prime}_{\rm H}$ is taken into account, these coefficients are consistent with those obtained in Refs. [Poisson:2004cw, Chatziioannou:2012gq].
Finally, for later convenience, we introduce the scalar tidal potentials [Poisson:2014gka]
$\mathscr{E}^{\rm q}\equiv\mathscr{E}_{ab}\,\Omega^{a}\Omega^{b},\qquad\mathscr{B}^{\rm q}\equiv\mathscr{B}_{ab}\,\Omega^{a}\Omega^{b},$ (4.12)
where $\Omega^{a}\equiv(\sqrt{1-z^{2}}\cos\psi,\,\sqrt{1-z^{2}}\sin\psi,\,z)$ is the unit spatial vector.
5 Secular dynamics in tidally deformed Kerr spacetime
A compelling and not yet fully explored question concerns how self-gravitating systems respond dynamically to external tidal fields. While tidal Love numbers have been extensively studied for compact objects in a wide range of theoretical frameworks [Chakraborty:2025wvs, Chia:2020yla, LeTiec:2020bos, LeTiec:2020spy, Pani:2015hfa, Gurlebeck:2015xpa, Poisson:2014gka, Kol:2011vg, Damour:2009va, Landry:2015zfa, Pani:2015nua, Binnington:2009bb, Chakrabarti:2013lua, Damour:2009vw, Chatziioannou:2020pqz, Flanagan:2007ix, Hinderer:2007mb, Cardoso:2017cfl, Pereniguez:2025jxq, Charalambous:2021mea, Perry:2023wmm], comparatively little attention has been devoted to the dynamics of matter orbiting a tidally distorted compact object.
Notable exceptions include studies of EMRI systems subject to external perturbations, which have shown modifications to orbital evolution and the possible onset of tidal resonances [Yang:2017aht, Yang:2019iqa, Bonga:2019ycj, Gupta:2021cno, Gupta:2022fbe].
Using the expression for the tidally deformed metric derived in Sec. 4, we investigate the motion of a test particle around a rotating black hole subject to an external tidal deformation. Our focus is on secular effects, namely corrections that accumulate over many orbital cycles, or equivalently on timescales much longer than the orbital period. As already mentioned in the Introduction, such effects are especially relevant for EMRIs, in which a small body orbits a much more massive black hole for a very large number of cycles.
In this regime, the dynamics is naturally characterized by a separation of timescales between the fast orbital motion and the much slower evolution induced by weak external perturbations. This hierarchy provides the basis for a description in terms of quantities averaged over the orbital period [Mino:2003yg, Hinderer:2008dm].
The presence of an external tidal field introduces additional long-term modifications to the orbital dynamics. Characterizing how such tidal deformations affect the secular behavior of orbital observables is therefore crucial, since even small corrections can accumulate over the inspiral and lead to potentially observable imprints in the phase of the emitted gravitational waves.
5.1 Secular Hamiltonian
We begin by examining the case of a test particle of mass $m$ moving on a timelike geodesic in the tidally deformed spacetime with metric $g_{\mu\nu}$. To describe the dynamics of this particle, we use the geodesic equation, including corrections to the Kerr geometry up to linear order in the metric perturbation $h_{\mu\nu}$. This approximation is valid provided that self-force effects can be neglected, namely in the limit $m/M\ll 1$, where $M$ denotes the mass of the Kerr black hole [Fujita:2016igj]. We also require $M/\mathscr{R}\ll 1$, where $\mathscr{R}$ is the length scale associated with the external tidal field, and we retain contributions only up to leading quadrupolar order, $\mathscr{O}\!\left((M/\mathscr{R})^{2}\right)$.
In the secular regime, and to first order in the metric perturbation $h_{\mu\nu}$, the physical orbit followed by the particle in the perturbed spacetime $g_{\mu\nu}=\tilde{g}_{\mu\nu}+h_{\mu\nu}$, which we denote by $\Gamma^{\prime}$, can be accurately approximated by an averaged geodesic $\Gamma$, defined within the same perturbed spacetime. This averaged geodesic can be interpreted as a secular orbit in the tidally deformed background.
For bound geodesic motion in Kerr spacetime, the radial, polar, and azimuthal motions are characterized by three independent fundamental frequencies. It is therefore convenient to parametrize the orbit using three phase variables $(q_{r},q_{z},q_{\psi})$, which label the oscillations associated with the radial, polar, and azimuthal motions, respectively, and evolve linearly with respect to a suitable parameter along the geodesic [Hinderer:2008dm, vandeMeent:2019cam, Schmidt:2002qk, Drasco:2003ky].
For a generic, non-resonant orbit, the secular average of a quantity $\mathscr{A}$ is defined as [Drasco:2005is]
$\langle\mathscr{A}\rangle=\frac{1}{(2\pi)^{3}}\int_{0}^{2\pi}dq_{\psi}\int_{0}^{2\pi}dq_{z}\int_{0}^{2\pi}dq_{r}\,\mathscr{A}(q_{r},q_{z},q_{\psi})\big|_{\Gamma}\,.$ (5.1)
As already mentioned, at the order considered here, the secularly averaged geodesic $\Gamma$ coincides with the physical orbit $\Gamma^{\prime}$. More precisely, differences arise only at higher order, for example when including first-order corrections in the mass ratio $m/M$ or second-order contributions in the metric perturbation $h_{\mu\nu}$ [Fujita:2016igj, Yang:2017aht, Camilloni:2023rra].
Moreover, to first order in the tidal perturbation, the secularly averaged energy $E$ and angular momentum $L$, both defined per unit rest mass, remain conserved along the secular orbit $\Gamma$ [Fujita:2016igj, Camilloni:2023rra].
The Hamiltonian of the test particle is given by
$H=\frac{1}{2}\,g_{\mu\nu}u^{\mu}u^{\nu}\,.$ (5.2)
Expanding to first order in the metric perturbation $h_{\mu\nu}$, we write
$H\simeq\frac{1}{2}\,\tilde{g}_{\mu\nu}\tilde{u}^{\mu}\tilde{u}^{\nu}+\tilde{g}_{\mu\nu}\tilde{u}^{\mu}u_{(1)}^{\nu}+\frac{1}{2}\,h_{\mu\nu}\tilde{u}^{\mu}\tilde{u}^{\nu}\,$, (5.3)
where we used that the four-velocity admits the expansion $u^{\mu}\simeq\tilde{u}^{\mu}+u_{(1)}^{\mu}$. Here, $\tilde{u}^{\mu}$ denotes the four-velocity along the geodesic of the unperturbed Kerr spacetime, while $u_{(1)}^{\mu}$ represents the correction induced by the tidal perturbation $h_{\mu\nu}$.
We now restrict our analysis to circular equatorial orbits ($z=0$). In this case, both the radial and polar motions are absent, and the only nonvanishing components of the four-velocity are $u_{v}$ and $u_{\psi}$, which can be expressed in terms of the conserved energy $E$ and angular momentum $L$.
Accordingly, the general secular average defined in Eq. (5.1) simplifies to an average over the azimuthal angle,
$\langle\mathscr{A}\rangle=\frac{1}{2\pi}\int_{0}^{2\pi}\mathscr{A}(\psi)\big|_{\Gamma}\,d\psi\,$, (5.4)
since the dependence on the radial and polar phases is absent, and the phase variable $q_{\psi}$ coincides with the azimuthal coordinate $\psi$ for circular equatorial motion.
As indicated in Eq. (3.2), the metric perturbation is given by twice the real part of the sum over modes with $m\in[-2,2]$. However, in the equatorial plane, after performing the secular average, only the $m=0$ component contributes, since the azimuthal averaging over a circular equatorial orbit eliminates all non-axisymmetric ($m\neq 0$) terms. Using the results of Sec. 4, we obtain [^3] This averaging does not affect the background metric $\tilde{g}_{\mu\nu}$, owing to the axial symmetry of the Kerr spacetime.
$\langle h_{vv}\rangle~=$ $~-\frac{\langle\mathscr{E}^{\rm q}\rangle}{r^{3}}\left(2a^{4}M-a^{2}r(2M^{2}+4Mr-3r^{2})+r^{3}(r-2M)^{2}\right)\,$, (5.5a)
$\langle h_{vz}\rangle~=$ $~a\frac{\langle\mathscr{B}^{\rm q}\rangle}{r}\left(a^{2}(r-2M)+2M^{2}r-r^{3}\right)\,$, (5.5b)
$\langle h_{v\psi}\rangle~=$ $~2a\frac{\langle\mathscr{E}^{\rm q}\rangle}{r^{3}}\left(a^{4}M-a^{2}r(M^{2}+Mr-r^{2})+r^{3}(M-r)^{2}\right)\,$, (5.5c)
$\langle h_{zz}\rangle~=$ $~\langle\mathscr{E}^{\rm q}\rangle r\left(a^{2}(2M-r)-2M^{2}r+r^{3}\right)\,$, (5.5d)
$\langle h_{z\psi}\rangle~=$ $~\frac{\langle\mathscr{B}^{\rm q}\rangle}{r}\left(a^{2}+r^{2}\right)\left(a^{2}(2M-r)-2M^{2}r+r^{3}\right)\,$, (5.5e)
$\langle h_{\psi\psi}\rangle~=$ $~-\frac{\langle\mathscr{E}^{\rm q}\rangle}{r^{3}}\left(2a^{6}M+a^{4}(r^{3}-2M^{2}r)-2a^{2}r^{4}(M-r)-2M^{2}r^{5}+r^{7}\right)\,$, (5.5f)
$\langle h_{vr}\rangle~=$ $~\langle h_{rr}\rangle~=\langle h_{rz}\rangle=\langle h_{r\phi}\rangle=0~,$ (5.5g)
where, using Eq. (4.12), the electric tidal scalar $\mathscr{E}^{q}$ is written explicitly as
$\mathscr{E}^{\rm q}$ $=\frac{\mathscr{E}_{11}-\mathscr{E}_{22}}{2}\left(1-z^{2}\right)\cos 2\psi+\frac{\mathscr{E}_{11}+\mathscr{E}_{22}}{2}\left(1-3z^{2}\right)+\mathscr{E}_{12}\left(1-z^{2}\right)\sin 2\psi$
$+2z\sqrt{1-z^{2}}\left(\mathscr{E}_{13}\cos\psi+\mathscr{E}_{23}\sin\psi\right)\,$, (5.6a)
and its secular average on the equatorial plane is given by $\langle\mathscr{E}^{q}\rangle\equiv(\mathscr{E}_{11}+\mathscr{E}_{22})/2$. Equivalently, one obtains an expression analogous to (5.6) for the magnetic tidal scalar $\mathscr{B}^{q}$, and its secular average on the equatorial plane is given by $\langle\mathscr{B}^{q}\rangle\equiv(\mathscr{B}_{11}+\mathscr{B}_{22})/2$.
We rescale $r\rightarrow r/M$, $a\rightarrow a/M$, and $L\rightarrow L/M$, so that these quantities are dimensionless. The conserved energy $E$ is already defined per unit rest mass and is therefore dimensionless.
Performing the secular average of the Hamiltonian to first order in the tidal perturbation yields
$\langle H\rangle=\langle H\rangle_{0}+\eta\,\langle H\rangle_{1}\,$, (5.7)
where
$\eta\equiv-\frac{M^{2}}{2}\,\langle\mathscr{E}^{\rm q}\rangle$ (5.8)
is a dimensionless perturbative parameter encoding information about the tidal environment [Yang:2017aht, Camilloni:2023rra, Grilli:2024fds]. We note that, although some components of the metric perturbation depend on the magnetic tidal scalar $\mathscr{B}^{\rm q}$, even after the averaging procedure, these components do not contribute to the Hamiltonian. As a result, any dependence on $\langle\mathscr{B}^{\rm q}\rangle$ drops out.
The unperturbed averaged Hamiltonian reads
$\langle H\rangle_{0}=-\frac{1}{2r\Delta}\left[\langle E^{2}\rangle\bigl(r^{3}+a^{2}(r^{2}+2)\bigr)+\langle L^{2}\rangle(r-2)-4a\langle E\rangle\langle L\rangle\right],$ (5.9)
while the first-order correction in $\eta$ is given by
$\langle H\rangle_{1}=\frac{1}{r^{3}}\left[\langle E^{2}\rangle\bigl(r^{5}+a^{2}r(3r^{2}-2)+2a^{4}\bigr)+\langle L^{2}\rangle\bigl(r^{3}-2r+2a^{2}\bigr)-4a\langle E\rangle\langle L\rangle\bigl(r^{3}-r+a^{2}\bigr)\right].$ (5.10)
In the zero-spin limit $a=0$, Eqs. (5.9) and (5.10) reproduce the results obtained in Refs. [Yang:2017aht, Camilloni:2023rra] for a tidally deformed Schwarzschild spacetime.
The secular Hamiltonian (5.7) provides a powerful and convenient framework to analyze long-term tidal effects on the location and properties of both the ISCO and the LR. These two orbits play a central role in black-hole physics, as they encode essential information about the behavior of matter, radiation, and gravitational perturbations in the strong-field region. In the next section, we examine in detail how tidal fields modify the properties of the ISCO and the LR.
6 Tidal deformation of ISCO and LR
In this section, we compute how the location and properties of the ISCO and of the LR of a Kerr black hole are modified by tidal effects, and we investigate how these modifications depend on the black-hole spin. As anticipated in Sec. 5, all results are obtained to first order in the tidal parameter $\eta$, using the secular Hamiltonian $\langle H\rangle$.
The ISCO is not only a mathematically well-defined geodesic of the Kerr geometry [Bardeen:1972fi], but also an object of direct astrophysical relevance. The binding energy at the ISCO sets the radiative efficiency of thin accretion disks, and measurements of disk spectra have therefore been used to infer black-hole spins through continuum-fitting methods [Narayan:2005ie, Zhang:1997dy]. Similarly, the $\mathrm{K}\alpha$ iron line originates in the innermost regions of the accretion disk, and its profile is sensitive to the location of the disk inner edge, which is often identified with the ISCO [Fabian:1989ej, Fabian:2000nu].
From a gravitational-wave perspective, the ISCO marks the end of the adiabatic inspiral and the onset of the plunge phase. In EMRI systems, even small tidal shifts in the location of the ISCO or in the associated orbital constants can accumulate over many cycles and lead to potentially detectable dephasings in the gravitational-wave signal [Ori:2000zn].
The LR is likewise of strong physical and observational relevance. Early studies showed that the LR controls the optical appearance of a collapsing star and the exponential fade-out of its observed luminosity [Ames:1968apj]. More recently, the LR has been recognized as playing a pivotal role in a variety of observational and theoretical contexts, including the determination of the photon-ring structure and, consequently, of the black-hole shadow [Falcke:1999pj]. The angular frequency and instability timescale (i.e., Lyapunov exponent) of the LR are directly related to the real and imaginary parts of black-hole quasinormal modes in the eikonal limit [Goebel:1972apj, Ferrari:1984zz, Mashhoon:1985cya, Berti:2005eb, Cardoso:2008bp, Cardoso:2021qqu, Yang:2012he]. Furthermore, the LR also determines the critical impact parameter relevant for high-energy scattering and threshold phenomena in numerical collision experiments [Pretorius:2007jn].
Computing tidal-induced secular shifts of LR quantities, such as the radius, orbital frequency, impact parameter, and instability rate, is therefore important, as such shifts could leave measurable imprints on black-hole shadows, ringdown signals, and critical scattering dynamics [Cardoso:2021qqu].
6.1 ISCO shifts
We begin by computing the shift induced by the tidal deformation on the location and properties of the ISCO of a Kerr black hole. We recall that, in the equatorial plane of the Kerr spacetime, there exist two ISCOs corresponding to prograde and retrograde orbits, respectively. These orbits can be determined by imposing the following three conditions [Isoyama:2014mja]
$\langle H\rangle\big|_{r=r_{\rm ISCO}}=-\frac{1}{2}\,,\qquad\frac{d\langle H\rangle}{dr}\bigg|_{r=r_{\rm ISCO}}=0\,,\qquad\frac{d^{2}\langle H\rangle}{dr^{2}}\bigg|_{r=r_{\rm ISCO}}=0\,.$ (6.1)
The solution of these equations yields the ISCO radius, as well as the specific energy and specific angular momentum of a test particle moving on this orbit. Up to first order in the perturbative parameter $\eta$, we write
$r^{\sigma}_{\rm ISCO}$ $\simeq r^{\sigma}_{0\,{\rm ISCO}}+\eta\,r^{\sigma}_{1\,{\rm ISCO}}\,$, (6.2)
$E^{\sigma}_{\rm ISCO}$ $\simeq E^{\sigma}_{0\,{\rm ISCO}}+\eta\,E^{\sigma}_{1\,{\rm ISCO}}\,$,
$L^{\sigma}_{\rm ISCO}$ $\simeq L^{\sigma}_{0\,{\rm ISCO}}+\eta\,L^{\sigma}_{1\,{\rm ISCO}}\,$,
where $\sigma=\pm 1$ labels prograde ($\sigma=+1$) and retrograde ($\sigma=-1$) orbits, respectively. Analogously, one can compute the orbital frequency $\Omega\equiv\langle u^{\psi}/u^{v}\rangle$ and the redshift invariant $U\equiv\langle u^{v}\rangle$ [Detweiler:2008ft, Sago:2008id, Shah:2012gu, vandeMeent:2015lxa], [^4] For circular equatorial orbits, these quantities coincide with $\langle u^{\phi}/u^{t}\rangle$ and $\langle u^{t}\rangle$ in Boyer-Lindquist coordinates, since $(v,\psi)$ differ from $(t,\phi)$ only by functions of the radial coordinate $r$ (cfr. Eq. (2.2)).
$\Omega^{\sigma}_{\rm ISCO}\simeq\Omega^{\sigma}_{0\,{\rm ISCO}}+\eta\,\Omega^{\sigma}_{1\,{\rm ISCO}}\,,\qquad U^{\sigma}_{\rm ISCO}\simeq U^{\sigma}_{0\,{\rm ISCO}}+\eta\,U^{\sigma}_{1\,{\rm ISCO}}\,.$ (6.3)
By solving Eqs. (6.1) at zeroth order in the tidal parameter $\eta$, one recovers the well-known expressions for the radial coordinate, the specific energy, and the specific angular momentum of the unperturbed ISCOs of a Kerr black hole, in agreement with the seminal results of Refs. [Chandrasekhar:1985kt, Bardeen:1972fi]:
$r^{\sigma}_{0\,{\rm ISCO}}=3+Z_{2}-\sigma\sqrt{(3-Z_{1})(3+Z_{1}+2Z_{2})}\,$, (6.4)
where
$Z_{1}=1+(1-a^{2})^{1/3}\left[(1-a)^{1/3}+(1+a)^{1/3}\right],\qquad Z_{2}=\sqrt{3a^{2}+Z_{1}^{2}}\,.$ (6.5)
The unperturbed energy and angular momentum of the ISCO can be conveniently expressed in terms of the unperturbed ISCO radius $r^{\sigma}_{0\,{\rm ISCO}}$, which, to simplify the notation, we denote from now on as $r_{0}$:
$E^{\sigma}_{0\,{\rm ISCO}}=\sqrt{1-\frac{2}{3r_{0}}}\,,\qquad L^{\sigma}_{0\,{\rm ISCO}}=a\sqrt{1-\frac{2}{3r_{0}}}+\sigma\,\frac{r_{0}}{\sqrt{3}}\,.$ (6.6)
Applying the same procedure to Eqs. (6.3) at leading order in $\eta$ yields
$\Omega^{\sigma}_{0\,{\rm ISCO}}=\frac{\sigma}{r_{0}^{3/2}+\sigma a}\,,\qquad U^{\sigma}_{0\,{\rm ISCO}}=\frac{r_{0}^{3/2}+\sigma a}{r_{0}^{3/4}\sqrt{r_{0}^{1/2}-(r_{0}^{3/2}-3)+2\sigma a}}\,.$ (6.7)
By solving Eqs. (6.1) to first order in the tidal parameter $\eta$, we can compute the leading perturbative corrections to the quantities characterizing the ISCO.
For the first-order tidal shift of the ISCO radius, we find
$r^{\sigma}_{1\,{\rm ISCO}}=-\frac{2\,\tilde{\Delta}(r_{0})}{r_{0}}\,\frac{\sqrt{r_{0}}\,\mathscr{A}(a,r_{0})+\sigma\,\mathscr{B}(a,r_{0})}{\mathscr{C}(a,r_{0})+\sigma\,\mathscr{D}(a,r_{0})}\,$, (6.8)
where
$\tilde{\Delta}(r_{0})\equiv\frac{\Delta(r_{0})}{M^{2}}=r_{0}^{2}-2r_{0}+a^{2}\,$, (6.9)
and the functions $\mathscr{A}$, $\mathscr{B}$, $\mathscr{C}$, and $\mathscr{D}$ are given by
$\mathscr{A}(a,r_{0})$ $=a^{6}(15r_{0}-84)+a^{4}(13r_{0}^{2}+30r_{0})+a^{2}(23r_{0}^{5}-106r_{0}^{4}+44r_{0}^{3}+72r_{0}^{2})$
$\quad+\bigl(4r_{0}^{8}-34r_{0}^{7}+101r_{0}^{6}-126r_{0}^{5}+48r_{0}^{4}\bigr)\,$, (6.10a)
$\mathscr{B}(a,r_{0})$ $=21a^{7}+a^{5}(93r_{0}-36r_{0}^{2})+a^{3}(3r_{0}^{5}+19r_{0}^{4}+9r_{0}^{3}-120r_{0}^{2})$
$\quad+a(19r_{0}^{7}-106r_{0}^{6}+200r_{0}^{5}-90r_{0}^{4}-12r_{0}^{3})\,$, (6.10b)
$\mathscr{C}(a,r_{0})$ $=a^{4}\sqrt{r_{0}}(34-9r_{0})-2a^{2}r_{0}^{3/2}\bigl[r_{0}(7r_{0}-48)+50\bigr]$
$\quad+r_{0}^{5/2}\bigl[(r_{0}-6)r_{0}(3r_{0}-16)-72\bigr]\,$, (6.10c)
$\mathscr{D}(a,r_{0})$ $=-9a^{5}-2a^{3}r_{0}(r_{0}+6)+ar_{0}^{2}\bigl[r_{0}(39r_{0}-172)+156\bigr]\,.$ (6.10d)
For the first-order tidal correction to the ISCO energy, we find
$E^{\sigma}_{1\,{\rm ISCO}}=\frac{\mathscr{P}_{E}(a,r_{0},r_{1})+\sigma\,\sqrt{r_{0}}\,\mathscr{Q}_{E}(a,r_{0},r_{1})}{2\,r_{0}^{11/4}\left[\sqrt{r_{0}}(r_{0}-3)+2a\sigma\right]^{3/2}}\,$, (6.11)
where, for notational convenience, we defined $r_{1}\equiv r^{\sigma}_{1\,{\rm ISCO}}$, and where the functions $\mathscr{P}_{E}$ and $\mathscr{Q}_{E}$ are given by
$\mathscr{P}_{E}(a,r_{0},r_{1})$ $=6a^{6}+6a^{4}r_{0}-4a^{4}r_{0}^{2}-a^{2}\left(8r_{0}^{5}-32r_{0}^{4}+22r_{0}^{3}+16r_{0}^{2}+3r_{0}r_{1}\right)$
$\quad-4r_{0}^{7}+22r_{0}^{6}-44r_{0}^{5}+32r_{0}^{4}+r_{0}^{3}r_{1}-6r_{0}^{2}r_{1}\,$, (6.12a)
$\mathscr{Q}_{E}(a,r_{0},r_{1})$ $=-16a^{5}-a^{3}\left(14r_{0}^{3}-22r_{0}^{2}-16r_{0}\right)$
$\quad-a\left(2r_{0}^{5}-14r_{0}^{4}+24r_{0}^{3}-4r_{0}^{2}-8r_{0}r_{1}\right)\,.$ (6.12b)
For the first-order tidal correction to the angular momentum, we find
$L^{\sigma}_{1\,{\rm ISCO}}=\frac{\mathscr{P}_{L}(a,r_{0},r_{1})+\sigma\,\sqrt{r_{0}}\,\mathscr{Q}_{L}(a,r_{0},r_{1})}{2\,r_{0}^{11/4}\left[\sqrt{r_{0}}(r_{0}-3)+2a\sigma\right]^{3/2}}\,$, (6.13)
where, as before, we defined $r_{1}\equiv r^{\sigma}_{1\,{\rm ISCO}}$, and the polynomials $\mathscr{P}_{L}$ and $\mathscr{Q}_{L}$ read
$\mathscr{P}_{L}(a,r_{0},r_{1})$ $=6a^{7}+6a^{5}r_{0}(1-2r_{0})-a^{3}r_{0}\left(6r_{0}^{4}-14r_{0}^{3}-18r_{0}^{2}+16r_{0}+3r_{1}\right)$
$\quad-ar_{0}^{2}\left(6r_{0}^{5}-36r_{0}^{4}+44r_{0}^{3}-4r_{0}^{2}-9r_{0}r_{1}+6r_{1}\right)\,$, (6.14a)
$\mathscr{Q}_{L}(a,r_{0},r_{1})$ $=2a^{6}(3r_{0}-8)-2a^{4}r_{0}\left(3r_{0}^{2}-8\right)$
$\quad-a^{2}r_{0}\left(2r_{0}^{5}+6r_{0}^{4}-22r_{0}^{3}+24r_{0}^{2}+3r_{0}r_{1}-4r_{0}-8r_{1}\right)$
$\quad-2r_{0}^{8}+10r_{0}^{7}-18r_{0}^{6}+4r_{0}^{5}+12r_{0}^{4}+r_{0}^{3}(r_{0}-6)r_{1}\,.$ (6.14b)
For the first-order tidal correction to the orbital angular frequency at the ISCO, we obtain
$\Omega^{\sigma}_{1\,{\rm ISCO}}=\frac{\mathscr{P}_{\Omega}(a,r_{0})+\sigma\,\mathscr{Q}_{\Omega}(a,r_{0},r_{1})}{2\sqrt{r_{0}}\,\left(r_{0}^{3/2}+\sigma a\right)^{2}}\,$, (6.15)
where the functions $\mathscr{P}_{\Omega}$ and $\mathscr{Q}_{\Omega}$ are given by
$\mathscr{P}_{\Omega}(a,r_{0})$ $=4a\sqrt{r_{0}}\,\left(r_{0}^{3}-a^{2}\right)\,$, (6.16a)
$\mathscr{Q}_{\Omega}(a,r_{0},r_{1})$ $=6a^{4}-2a^{2}r_{0}(r_{0}+3)-2r_{0}^{5}+4r_{0}^{2}-3r_{0}r_{1}\,.$ (6.16b)
Finally, the first-order tidal correction to the redshift invariant reads
$U^{\sigma}_{1\,{\rm ISCO}}=\frac{\mathscr{P}_{U}(a,r_{0},r_{1})+2\sigma\sqrt{r_{0}}\,\mathscr{Q}_{U}(a,r_{0},r_{1})}{2r_{0}^{11/4}\left[\sqrt{r_{0}}(r_{0}-3)+2\sigma a\right]^{3/2}}\,$, (6.17)
where
$\mathscr{P}_{U}(a,r_{0},r_{1})$ $=6a^{6}+6a^{4}r_{0}(2r_{0}-1)+a^{2}r_{0}\left(4r_{0}^{4}-24r_{0}^{3}-6r_{0}^{2}+12r_{0}-3r_{1}\right)$
$\quad-r_{0}^{3}\left(6r_{0}^{3}-8r_{0}^{2}+3r_{1}\right)\,$, (6.18a)
$\mathscr{Q}_{U}(a,r_{0},r_{1})$ $=-6a^{5}+a^{3}r_{0}\left(7r_{0}^{2}-9r_{0}+6\right)+ar_{0}\left(3r_{0}^{4}-3r_{0}^{3}+8r_{0}^{2}-6r_{0}+3r_{1}\right)\,.$ (6.18b)
These expressions reveal a strong dependence of the tidal ISCO shifts on both the black-hole spin and the relative orientation of the orbit. In Fig. 1 we illustrate the dependence of these quantities on the spin of the Kerr black hole. We comment further on these results below. In the non-spinning limit $a=0$, all tidal corrections smoothly reduce to those computed for a tidally deformed Schwarzschild black hole in Ref. [Camilloni:2023rra].
6.2 LR shifts
Similarly, one can analyze how the tidal environment affects the location and properties of the LR of a Kerr black hole by solving the following conditions in terms of the secular Hamiltonian:
$\langle H\rangle\big|_{r=r_{\rm LR}}=0\,,\qquad\frac{d\langle H\rangle}{dr}\bigg|_{r=r_{\rm LR}}=0\,.$ (6.19)
The solution of these equations yields the LR radius, the impact parameter, and the orbital frequency, which completely characterize the LR. We define the impact parameter as
$b\equiv\sigma\,\frac{L}{E},$
where $\sigma=\pm 1$ labels prograde and retrograde orbits, respectively. With this convention, the conserved angular momentum $L$ is positive (negative) for prograde (retrograde) motion, so that $b$ is positive for both orientations and retains its standard geometric interpretation as a distance.
Up to first order in the perturbative parameter $\eta$, we write
$r^{\sigma}_{\rm LR}$ $\simeq r^{\sigma}_{0\,{\rm LR}}+\eta\,r^{\sigma}_{1\,{\rm LR}}\,$, (6.20)
$b^{\sigma}_{\rm LR}$ $\simeq b^{\sigma}_{0\,{\rm LR}}+\eta\,b^{\sigma}_{1\,{\rm LR}}\,$,
$\Omega^{\sigma}_{\rm LR}$ $\simeq\Omega^{\sigma}_{0\,{\rm LR}}+\eta\,\Omega^{\sigma}_{1\,{\rm LR}}\,.$
By solving Eqs. (6.19) to zeroth order in the tidal parameter $\eta$, we recover the known results for the radial coordinate, impact parameter, and frequency of the unperturbed LRs of a Kerr black hole [Chandrasekhar:1985kt, Bardeen:1972fi]
$r^{\sigma}_{0\,{\rm{LR}}}=2\left(1+\cos\left[\frac{2}{3}\cos^{-1}(-\sigma a)\right]\right)\,,~~b^{\sigma}_{0\,{\rm{LR}}}=\frac{\hat{r}_{0}^{2}-2\sigma a\hat{r}_{0}^{1/2}+a^{2}}{\hat{r}_{0}^{3/2}-2\hat{r}_{0}^{1/2}+\sigma a}\,,~~\Omega^{\sigma}_{0\,{\rm{LR}}}=\frac{\sigma}{\hat{r}^{3/2}_{0}+\sigma a}\,$ (6.21)
where, to simplify the notation, we defined the unperturbed LR radius $\hat{r}_{0}\equiv r^{\sigma}_{0\,{\rm{LR}}}$. Here, the LR angular frequency is defined analogously as the secularly averaged quantity $\Omega\equiv\langle u^{\psi}/u^{v}\rangle$, evaluated on circular equatorial null orbits.
By solving Eqs. (6.19) to first order in the tidal parameter $\eta$, we obtain the leading tidal corrections to the LR quantities.
For the first-order shift of the LR radius, we find
$r^{\sigma}_{1\,{\rm LR}}=\frac{2\left[\mathscr{F}(a,\hat{r}_{0})+\sigma\sqrt{\hat{r}_{0}}\,\mathscr{G}(a,\hat{r}_{0})\right]}{\hat{r}_{0}\left(\hat{r}_{0}^{2}-6\hat{r}_{0}-3a^{2}+8\sigma a\sqrt{\hat{r}_{0}}\right)}\,$, (6.22)
where the functions $\mathscr{F}$ and $\mathscr{G}$ are given by
$\mathscr{F}(a,\hat{r}_{0})$ $=-3a^{6}+a^{4}(2\hat{r}_{0}^{2}-3\hat{r}_{0})+a^{2}(4\hat{r}_{0}^{5}-16\hat{r}_{0}^{4}+11\hat{r}_{0}^{3}+8\hat{r}_{0}^{2})$
$\quad+\left(2\hat{r}_{0}^{7}-11\hat{r}_{0}^{6}+22\hat{r}_{0}^{5}-16\hat{r}_{0}^{4}\right)\,$, (6.23a)
$\mathscr{G}(a,\hat{r}_{0})$ $=8a^{5}+a^{3}(7\hat{r}_{0}^{3}-11\hat{r}_{0}^{2}-8\hat{r}_{0})+a\left(\hat{r}_{0}^{5}-7\hat{r}_{0}^{4}+12\hat{r}_{0}^{3}-2\hat{r}_{0}^{2}\right)\,.$ (6.23b)
For the first-order tidal correction to the impact parameter, we obtain
$b^{\sigma}_{1\,{\rm LR}}=\frac{\mathscr{R}_{b}(a,\hat{r}_{0})+\sigma\sqrt{\hat{r}_{0}}\,\mathscr{K}_{b}(a,\hat{r}_{0})}{(\hat{r}_{0}-2)\sqrt{\hat{r}_{0}}+\sigma a}\,$, (6.24)
where the functions $\mathscr{R}_{b}$ and $\mathscr{K}_{b}$ read
$\mathscr{R}_{b}(a,\hat{r}_{0})$ $=2a^{4}+a^{2}(3\hat{r}_{0}^{3}-5\hat{r}_{0}^{2})+\left(\hat{r}_{0}^{5}-3\hat{r}_{0}^{4}+4\hat{r}_{0}^{3}-2\hat{r}_{0}^{2}\right)\,$, (6.25a)
$\mathscr{K}_{b}(a,\hat{r}_{0})$ $=2a^{3}(\hat{r}_{0}-2)-2a\,\hat{r}_{0}(\hat{r}_{0}^{2}-2)\,.$ (6.25b)
Finally, we find that the angular frequency of the LR receives a tidal correction given by
$\Omega^{\sigma}_{1\,{\rm LR}}=\frac{\mathscr{R}_{\Omega}(a,\hat{r}_{0},\hat{r}_{1})+\sigma\,\sqrt{\hat{r}_{0}}\,\mathscr{K}_{\Omega}(a,\hat{r}_{0},\hat{r}_{1})}{\tilde{\Delta}(\hat{r}_{0})\,\left(\hat{r}_{0}^{3/2}+\sigma a\right)^{2}}\,$, (6.26)
where $\hat{r}_{1}\equiv r^{\sigma}_{1\,{\rm LR}}$, and the functions $\mathscr{R}_{\Omega}$ and $\mathscr{K}_{\Omega}$ read
$\mathscr{R}_{\Omega}(a,\hat{r}_{0},\hat{r}_{1})$ $=6a^{5}+a^{3}\left(9\hat{r}_{0}^{3}-13\hat{r}_{0}^{2}-4\hat{r}_{0}\right)$
$\quad+a\left(3\hat{r}_{0}^{5}-11\hat{r}_{0}^{4}+12\hat{r}_{0}^{3}-2\hat{r}_{0}^{2}-4\hat{r}_{0}\hat{r}_{1}\right)\,$, (6.27a)
$\mathscr{K}_{\Omega}(a,\hat{r}_{0},\hat{r}_{1})$ $=4a^{4}(\hat{r}_{0}-3)+a^{2}\left(3\hat{r}_{0}^{4}-17\hat{r}_{0}^{3}+10\hat{r}_{0}^{2}+16\hat{r}_{0}\right)$
$\quad+\hat{r}_{0}^{6}-9\hat{r}_{0}^{5}+22\hat{r}_{0}^{4}-14\hat{r}_{0}^{3}-4\hat{r}_{0}^{2}-2\hat{r}_{0}^{2}\hat{r}_{1}+6\hat{r}_{0}\hat{r}_{1}\,.$ (6.27b)
Note that, in the non-spinning limit $a=0$, this expression correctly reduces to the first-order tidal correction to the light-ring frequency of a Schwarzschild black hole derived in Ref. [Camilloni:2023rra].
The analytical results obtained for the tidal deformations of the ISCO and the LR are derived for a generic external tidal environment. The only assumptions are that the characteristic length scale of the Kerr black hole, set by its mass $M$, is much smaller than the curvature radius $\mathscr{R}$ of the external spacetime ($M\ll\mathscr{R}$), and that the tidal perturbation is retained only up to the leading quadrupolar order.
We illustrate these results in Figs. 1 and 2. In both cases, the plots reveal a pronounced dependence of the tidal corrections to the ISCO and the LR on the black-hole spin parameter $a$.
Specifically, the left panels of both figures show that, for prograde orbits, tidal shifts are increasingly suppressed as the black-hole spin grows, relative to the corresponding non-spinning case. In contrast, the right panels show that for retrograde orbits the trend is reversed: increasing the black-hole spin leads to progressively enhanced tidal deformations.
This behavior is consistent across all orbital quantities considered for both the ISCO and the LR, and admits a clear physical interpretation. As originally discussed by Bardeen and collaborators [Bardeen:1972fi], increasing the black-hole spin causes prograde circular orbits to move closer to the event horizon $r_{+}$, while retrograde orbits are pushed to larger radii. As a consequence, prograde ISCO and LR orbits probe regions of stronger gravitational binding, where the coupling to a fixed external tidal field is effectively reduced, leading to a suppression of tidal effects as $a$ increases. Conversely, retrograde orbits reside at larger radii for higher spins, where the tidal interaction is stronger, resulting in enhanced tidal deformations.
Finally, while several analyses in the literature restrict attention to the subextremal regime $0<|a|<M$, we find no indication of pathological behavior as $|a|\rightarrow M$ in the static ($\omega=0$) sector considered here. In particular, the metric perturbation reconstructed in the ORG, as well as the associated observable quantities derived from the secular Hamiltonian, remain regular in this limit. This suggests that, within the assumptions of our framework, namely vacuum, quadrupolar, and slowly varying tidal perturbations, the extremal Kerr case is approached smoothly.
[…]
Figure 1: Secular tidal perturbations to the radius $r$, energy $E$, angular momentum $L$, frequency $\Omega$, and redshift factor $U$ of a massive test particle moving on the ISCO of a tidally deformed Kerr black hole. The shifts are shown as functions of the spin $a$ and are normalized to the Schwarzschild ($a=0$) case. The left panel corresponds to prograde motion, while the right panel corresponds to retrograde motion.
[…]
Figure 2: Secular tidal perturbations to the radius $r$, impact parameter $b$, and frequency $\Omega$ of a massless test particle moving on the LR of a tidally deformed Kerr black hole. The shifts are shown as functions of the spin $a$ and are normalized to the Schwarzschild ($a=0$) case. The left panel corresponds to prograde motion, while the right panel corresponds to retrograde motion.
7 Conclusions
In this paper, we developed an analytic framework to describe the tidal deformation of rotating black holes and applied it to the study of the secular dynamics of test particles in the strong-field regime. By solving the Teukolsky master equation in the Hartle–Hawking tetrad for static modes and employing metric reconstruction techniques in the outgoing radiation gauge, we derived fully explicit expressions for the quadrupolar metric perturbation of a Kerr black hole subject to slowly varying external tidal fields. In the non-spinning limit, our results smoothly reduce to the known expressions for tidally deformed Schwarzschild black holes [Binnington:2009bb]. All results are obtained in the near-zone regime $M\ll\mathscr{R}$ and at leading quadrupolar order in the tidal deformation.
Using the reconstructed metric, we investigated the secular dynamics of test particles in the tidally deformed Kerr spacetime and derived the corresponding secular Hamiltonian to first order in the tidal perturbation. This allowed us to compute tidal-induced shifts in the location and properties of the innermost stable circular orbit and the light ring. We found that these shifts exhibit a strong dependence on the black-hole spin and on the relative orientation between the orbital angular momentum and the spin. In particular, tidal effects are suppressed for prograde orbits around rapidly rotating black holes, while they are significantly enhanced for retrograde orbits. These results highlight the role of spin-tidal couplings in shaping the strong-field orbital dynamics of matter and radiation in the near-horizon region, and extend to spinning black holes the strong-field tidal effects previously identified in charged black-hole spacetimes [Grilli:2024fds].
Our results represent a step toward a more complete understanding of the relativistic tidal response of spinning black holes. In particular, the fully analytic nature of our framework provides controlled access to spin–tidal couplings and secular strong-field effects in rotating black-hole spacetimes, bridging perturbative tidal analyses and applications relevant to gravitational-wave observations.
Several directions for future work emerge from this analysis. An important extension is the inclusion of higher-order tidal multipoles, as well as the study of time-dependent tidal perturbations. It would also be relevant to compute the associated energy and angular-momentum fluxes, following existing approaches in black-hole perturbation theory [Poisson:2004cw, Chatziioannou:2012gq, Berens:2024czo]. These extensions are likely to be particularly relevant for the modeling of EMRIs and for future tests of general relativity with space-based gravitational-wave detectors such as LISA.
A further extension is to move beyond the circular and equatorial approximation adopted in this paper. The present framework can be extended to generic bound geodesics in a tidally deformed Kerr spacetime, including eccentric and inclined orbits, by employing the general secular-averaging procedure given in Eq. (5.1). This would allow one to investigate how external tidal fields influence orbital precession, resonances, and long-term secular evolution in EMRI systems beyond the circular limit.
A related but distinct application concerns hierarchical triple systems, in which the external tidal field is generated by a third massive black hole. Recent work has shown that tidal interactions in such systems can significantly affect the eccentricity evolution and merger timescale of compact binaries treated within post-Newtonian theory [Camilloni:2023xvf, Cocco:2025adu, Cocco:2025udb]. The tidally deformed Kerr metric derived here provides a strong-field framework to study analogous effects when the inner object is modeled as a test particle orbiting a spinning black hole, thereby enabling a systematic comparison between EMRI and comparable-mass regimes. We plan to explore these extensions in future work.
Acknowledgements
M. Cocco, G. Grignani, and M. Orselli acknowledge financial support from the Italian Ministry of University and Research (MUR) through the program “Dipartimenti di Eccellenza 2018–2022” (Grant SUPER-C). G. Grignani and M. Orselli also acknowledge support from the “Fondo di Ricerca d’Ateneo” 2023 (GraMB) of the University of Perugia, and from the Italian Ministry of University and Research (MUR) via PRIN 2022ZHYFA2, GRavitational wavEform models for coalescing compAct binaries with eccenTricity (GREAT).
M. Cocco, T. Harmark, M. van de Meent, and M. Orselli acknowledge support from the “Center of Gravity”, a Center of Excellence funded by the Danish National Research Foundation under Grant No. DNRF184. M. van de Meent further acknowledges financial support from the VILLUM Foundation (Grant No. VIL37766), the DNRF Chair program (Grant No. DNRF162) funded by the Danish National Research Foundation, and the European Union’s Horizon ERC Synergy Grant Making Sense of the Unexpected in the Gravitational-Wave Sky (GWSky–101167314).